### v1

# Sequential Conformal Risk Control UQ Uncertainty Quantification

Abstract
Recent advances in object detectors have led to their adoption for industrial uses. However, their deployment in critical applications is hindered by the inherent lack of reliability of neural networks and the complex structure of object detection models. To address these challenges, we turn to Conformal Prediction, a post-hoc procedure which offers statistical guarantees that are valid for any dataset size, without requiring prior knowledge on the model or data distribution. Our contribution is manifold: first, we formally define the problem of Conformal Object Detection (COD) and introduce a novel method, Sequential Conformal Risk Control (SeqCRC), that extends the statistical guarantees of Conformal Risk Control (CRC) to two sequential tasks with two parameters, as required in the COD setting. Then, we propose loss functions and prediction sets suited to applying CRC to different applications and certification requirements. Finally, we present a conformal toolkit, enabling replication and further exploration of our methods. Using this toolkit, we perform extensive experiments, yielding a benchmark that validates the investigated methods and emphasizes trade-offs and other practical consequences.
Index Terms:
Conformal Prediction, Conformal Risk Control, Object Detection, AI Safety, Uncertainty Quantification
Figure 1: Conformal Object Detection Pipeline. Our method corrects the output of a given object detector $f$ with finite-sample probabilistic guarantees at a desired error rate $\alpha$. The figure illustrates the key stages: (1) Red bounding boxes show the initial raw predictions from the object detector $f(X_{i})$ which are then filtered by Non-Maximum Suppression to a smaller set. (2) Conformal Confidence Thresholding applies a statistically grounded confidence threshold, resulting in selected predictions shown with yellow bounding boxes with a guaranteed risk. (3) Conformal Localization corrects the previous dashed yellow boxes to build the final statistically valid purple bounding boxes. For each localized object, Conformal Classification constructs a prediction set of labels (e.g., "Set: {Boat, Plane}") from the model’s predictions, guaranteed to contain the true class with high probability. These steps together control a user-defined risk in localization and classification at a predetermined level $\alpha^{\mathrm{tot}}$.
I Introduction
Object Detection (OD) has emerged as a pivotal task in computer vision, aiming to localize and classify objects within images or video sequences. It has had a remarkable impact in diverse fields, ranging from autonomous vehicles [1] to medical imaging [2]. However, there is an increasing demand to certify predictions of OD models within critical systems [3, 4, 5, 6].
In this light, while existing Uncertainty Quantification (UQ) methods are generally Bayesian [7, 8, 9] or heuristic [10, 11, 12], our approach emphasizes distribution-free UQ methods [13]. These methods neither assume prior knowledge of the data distribution nor are restricted to a particular class of predictive models. They are statistically valid even with a finite number of examples (i.e. non-asymptotic). When combining object detection with these methodologies, most notably Conformal Prediction (CP), we take an a posteriori strategy. This post-hoc approach is theoretically sound and computationally lightweight, allowing for rapid integration into real-world systems. Moreover, while the existing literature on conformal prediction for OD [14, 15, 16] and tolerance regions for OD [17, 18] focuses on specific steps of the whole OD pipeline or is sometimes tailored to specific models or tasks, our approach addresses a more unifying framework. It is applicable to virtually any object detectors, and readily deployable in existing systems, ensuring reliability without a comprehensive overhaul of established frameworks, and at very little cost.
After discussing related works and formally defining the setup (Section II), we make the following set of contributions:
  1. 1.
    In Section III, we introduce a novel approach called Sequential Conformal Risk Control (SeqCRC), which enables control over two task-specific losses parameterized by two sequentially selected parameters—a key requirement for general OD tasks. This allows the user to reliably tune a confidence threshold and to compute reliable “error margins” both for the localization and classification tasks.
  2. 2.
    In Section IV, we review several state-of-the-art conformal loss functions and introduce new ones, discussing their relevance and the trade-offs they entail. We also examine the crucial choice of metric for matching predictions with ground truths.
  3. 3.
    In Section V, we empirically validate the correctness, efficiency, and broad applicability of our framework, via the first large benchmark of Conformal OD in the literature.
  4. 4.
    Finally, we provide a comprehensive, open-source codebase111 https://github.com/leoandeol/cods—a repository of models and conformal procedures—that facilitates the development of new guarantees, loss functions, and methodologies.
I-A Related Works
We now discuss several approaches to UQ for OD, with or without statistical guarantees. The first set of references below is closest to our paper.
Conformal prediction for OD. Conformal prediction [13, 19] and variants [20, 21, 22] are natural post-processing methods that offer statistically valid uncertainty estimates. Some applications to object detection have been proposed recently, notably through conformal prediction in [14, 16, 15, 23] and tolerance regions in [17, 18]. These two frameworks build and calibrate post-hoc prediction sets with guarantees, although slightly different ones. Most works have focused, however, only on the localization task (i.e., a regression task on the coordinates of a bounding box on the image). Two exceptions are [23], and [17]. The former, includes a guarantee on the classification task, while also producing conformal corrections in localization conditional on the class of the predicted object. The latter, builds all-encompassing sets, including confidence, classification and localization that are specifically tailored to Faster R-CNN models. However, none offers a holistic framework to build sets. That is, prediction sets that would be valid for virtually all families of OD models with simultaneous statistical guarantees on the number of object, as well as their localization, and classification.
A similar theoretical approach to ours, with two parameters to tune sequentially, was recently and independently introduced in [24] for rank retrieval. However, their theoretical guarantee is either asymptotic or requires two data splits for finite samples, whereas our method provides a finite-sample guarantee using a single data split only. The framework of LTT [25] provides another method to build prediction sets, which can depend on more than one parameter, using a multiple testing approach. This framework has recently been applied to Rank Retrieval [24] and to Language Modeling [26]. It has also been applied to OD to simultaneously control precision and recall in a medical setting in [18].
Other approaches to UQ for OD. Leaving statistical guarantees aside, let us mention other useful methods to compute uncertainty estimates for OD. First, several heuristic approaches exist. MetaDetect [10] trains an auxiliary network on top of the object detector to estimate predictive uncertainty from handpicked metrics. It does so by learning to discriminate false positives from true positives (in terms of Intersection over Union, IoU), and to predict IoU. The method of [27] allows to calibrate the confidence score of predicted bounded boxes, in a position-aware manner. The resulting confidence of bounding boxes in all regions of the image is then improved in terms of calibration error. Estimating uncertainty from the gradient of the loss (with regards to the the weights) has also been studied in [28] and a comparison of the two approaches is discussed in [29]. Notably, both approaches offer significantly more insights into a bounding box’s likelihood of containing an object than the native confidence score.
Another widely used approach is Bayesian inference, which models uncertainty through a posterior distribution over parameters. This distribution cannot be computed analytically and needs to be approximated. It is mainly done using methods such as Monte-Carlo Dropout [30, 31] or Variational Inference [32, 33], but alternatives exist [34, 35]. A first application of Bayesian Neural Networks to OD [7] demonstrated significantly increased precision and recall, and a good estimation label uncertainty in the open set detection setting. Another approach named BayesOD [8], applicable to anchor-based networks, allows to replace NMS with a Bayesian fusion: after clustering largely overlapping predictions, the clusters are merged into a single prediction with an uncertainty estimate on its coordinates. This allows incorporating information from less confident overlapping predictions, rather than simply deleting them. Other, non Bayesian improvements to the NMS have also been proposed [36, 37].
All these works are complementary to ours. Indeed, even if they lack theoretical guarantees, they provide uncertainty estimates that could be used as inputs to our method, in order to construct more adaptive prediction sets. Our approach enables the integration of standard uncertainty metrics to adjust predictions in an uncertainty-aware manner, for example, by substantially enlarging prediction sets for highly uncertain cases, while leaving confident ones unchanged, instead of applying a uniform scaling factor.
II Setting: Object Detection with Uncertainty Quantification
Both Object Detection (OD) and Uncertainty Quantification (UQ) for OD have long been dominated by heuristic methods. As we aim to bring guaranteed uncertainty quantification to this field, we formally introduce the setting, the problem considered, and previous works. Our new SeqCRC method and theoretical analysis are presented in Section III, while a large selection of OD-specific modeling choices are discussed in Section IV.
II-A Problem Setup (without UQ)
Figure 2: Abstract diagram representing the lens through which we study object detectors. As seen, we consider the NMS to have already been conducted post-prediction.
The task of object detection is complex, as it involves both localizing and classifying, in an image, an a priori unknown number of objects. To that end, most models (e.g. [38, 39, 40, 41]) predict a large amount of bounding boxes with an additional confidence component. This score is further used to filter out most predicted boxes using a threshold to keep the most relevant bounding boxes. While existing models compute this score in different ways, and some may use additional scores [42], we only consider a single score, irrespective of its construction, to provide a holistic setting and method.
Formal Problem Setup. We denote by $\mathcal{X}$ the input (image) space. Each object on an image is modeled as an element of ×$\mathcal{B}\times\mathcal{C}$, where =+$\mathcal{B}=\mathbb{R}^{4}_{+}$ is the bounding box space and $\mathcal{C}=\left\{1,\cdots,K\right\}$ is the label space ($K$ is the number of labels). Each location vector $b=(b^{\leftarrow},b^{\uparrow},b^{\rightarrow},b^{\downarrow})\in\mathcal{B}$ contains the coordinates of the top-left and bottom-right corners of its corresponding bounding box, respectively $(b^{\leftarrow},b^{\uparrow})$ and $(b^{\rightarrow},b^{\downarrow})$. We assume that $b^{\leftarrow}\leq b^{\rightarrow}$ and $b^{\uparrow}\leq b^{\downarrow}$. [^2] The horizontal axis points to the right while the vertical axis heads downwards.
Throughout this paper, we will refer to inclusion of a bounding box $b$ by another $\tilde{b}$, denoted $b\subseteq\tilde{b}$, to mean that we simultaneously have $b^{\leftarrow}\geq\tilde{b}^{\leftarrow},b^{\uparrow}\geq\tilde{b}^{\uparrow},b^{\rightarrow}\leq\tilde{b}^{\rightarrow},b^{\downarrow}\leq\tilde{b}^{\downarrow}$. We use the notation $b\not\subseteq\tilde{b}$ whenever one or more of these inequalities do not hold true.
Commonly in OD, the output for a predicted object is not a single label, but an estimated probability distribution on $\mathcal{C}=\{1,\cdots,K\}$, such as a softmax output. Further recall that we work with a predictor $f$ that outputs a large, fixed number $N^{\text{pred}}$ of objects. Those objects are each composed of a bounding box, with an additional classification score (softmax output) and a confidence score in $[0,1]$. Such predictors are therefore functions $f:\mathcal{X}\rightarrow\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\text{pred}}}$, where $\Sigma_{K-1}$ denotes the $(K-1)$-simplex. [^3] That is, $\Sigma_{K-1}=\{c\in[0,1]^{K}:\sum_{i=1}^{K}c_{i}=1$}.
Importantly, a recurring post-processing step for object detectors is Non-Maximum Suppression (NMS), which consists in removing redundancies and thus significantly lowering the amount of predicted boxes. Its output lies in the reduced space (×$\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\mathrm{nms}}_{i}}$ where $N^{\mathrm{nms}}_{i}\ll N^{\text{pred}}$ and varies for each image $X_{i}$.
Instances of the Problem. Most of the commonly used models in the literature fit in the setting presented above. Popular examples are YOLO [38, 39, 40] for single-stage methods, and Faster R-CNN [42] for two-stage methods. Other examples include the more recent transformer-based models such as DETR [43] and its variants [44, 45], as well as the diffusion-inspired model DiffusionDet [46].
II-B Adding UQ with three parameters
In this work, we study an approach to UQ in OD that has gained renewed interest over the past decade: Conformal Prediction [19]. We consider a post-processing approach that corrects predictions, such as adding a margin on a bounding box or outputting multiple classes instead of the most likely one. This is the conformal approach, and such corrected bounding boxes, or sets of classes, are called “prediction sets”. Section I-A discusses earlier works following this approach. Put briefly, our conformal prediction method for OD is a simple, generic, and statistically-grounded way to tune three parameters:
  • a confidence parameter $\lambda^{\mathrm{cnf}}\in\Lambda^{\mathrm{cnf}}$ used to define a threshold below which predicted objects are ignored;
  • a localization parameter $\lambda^{\mathrm{loc}}\in\Lambda^{\mathrm{loc}}$ used to correct the size of predicted bounding boxes (by, e.g., adding a margin in pixels to these), and
  • a classification parameter $\lambda^{\mathrm{cls}}\in\Lambda^{\mathrm{cls}}$ used to build the set of most likely classes for any predicted object (via, e.g., a threshold on the softmax scores).
Figure 1 illustrates the usage of these three parameters. After Non-Maximum Suppression (which is considered as part of the OD model itself), the thresholding step (using $\lambda^{\mathrm{cnf}}$) allows us to reduce the list of predicted objects to a smaller list of objects $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$ that are more likely to correspond to true objects. Then, from these predictions, we carry out two tasks in parallel. For localization, we construct a list $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ of bounding boxes in $\mathcal{B}$ obtained by expanding the bounding boxes from $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$ (with an additive or multiplicative margin). For classification, we construct a list $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ of subsets of $\{1,\ldots,K\}$—one subset for each predicted object, with the labels predicted as most likely.
In the rest of the paper, we show how to construct the prediction sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}},\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}},\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ and how to tune the parameters $\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}},\lambda^{\mathrm{cls}}$ so that the resulting predictions are statistically valid (Section III) and practically meaningful (Section IV). A comprehensive list of experiments is discussed in Section V.
III Conformal Object Detection
Neural networks are often regarded as black boxes due to their complexity. There is an important research effort in deep learning theory for generalization bounds [47, 48, 49]. However, due to the complexity of neural networks, these bounds are often loose and therefore not yet suitable for industrial requirements. In particular, in Object Detection, there is a lack of a theoretical framework. This section introduces a post-hoc method for constructing model-agnostic, distribution-free prediction sets guaranteed to contain the ground truth with large probability. This method pertains to an approach known as Conformal Prediction and first introduced by Vovk et al. (e.g., [13] and references therein). In particular, one algorithmic variant called Split (or Inductive) Conformal Prediction [50] is well suited to post-hoc applications on large models. This simple approach allows us to bypass generalization error bounds of neural networks, via a simple yet statistically rigorous “testing” viewpoint. For an introduction to the subject we refer the reader to, e.g., [19, 51, 13, 52].
In this paper, we study an extension of Split Conformal Prediction called Conformal Risk Control (CRC) [53] and that adds flexibility for the user—a necessary feature for many critical applications. Next we start with some useful background on CRC, which only applies to prediction sets indexed by a single (scalar) parameter. Then, in Section III-B, we design and analyze an extension of CRC suited to our Object Detection setting, with three dependent scalar parameters.
III-A Background: Conformal Risk Control
In this section we recall how conformal risk control proceeds in the post-hoc setting, [^4] By post-hoc we refer to a procedure conducted on top of a pretrained model’s predictions. for a general supervised learning task such as classification or regression. Assume that we are given a deterministic predictor $f$ [^5] The predictor is in fact required to be independent from the test and the calibration data. Roughly speaking, this corresponds to a setting where the model $f$ is trained and validated on a different split or pretrained. To construct a prediction set, we require a calibration dataset $\mathcal{D}_{\text{cal}}=(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ of $n$ examples and a test point $X_{\mathrm{test}}\in\mathcal{X}$. We then build a prediction set $\Gamma_{\lambda}(X_{\mathrm{test}})$666 Equivalently, prediction sets can be expressed in terms of a non-conformity score. See [54] for details. with the goal of controlling a loss $\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$ below a user-defined threshold $\alpha$. For example, in multiclass classification where $\mathcal{Y}=\{1,\ldots,K\}$, we can work with $\Gamma_{\lambda}(X_{\mathrm{test}})\subset\mathcal{Y}$ and $\ell(y,S)=\mathbbm{1}_{y\notin S}$. Multiple other examples exist, with non-binary losses or prediction sets $\Gamma_{\lambda}(X_{\mathrm{test}})$ that are not necessarily subsets of $\mathcal{Y}$ as, e.g., for the localization task (in that case, the prediction set is the predicted box after application of a margin, see Section IV-C). For notational simplicity, we denote the random losses777 These are referred to as random losses because they are functions of random variables, hence random variables themselves. by $L_{i}(\lambda)=\ell(Y_{i},\Gamma_{\lambda}(X_{i}))$, for $i=1,...,n$. We also denote the random loss of the test sample by $L_{\mathrm{test}}(\lambda)=\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$.
Typically, such prediction sets $\Gamma_{\lambda}(x)$ are parameterized by a real number $\lambda$ that controls their sizes. Conformal methods are data-driven procedures for computing some $\widehat{\lambda}$ whose associated prediction sets $\Gamma_{\widehat{\lambda}}(X_{\mathrm{test}})$ are both statistically valid and as “small” as possible. (As will become clear later, $\widehat{\lambda}$ depends on the calibration data, and thus $\Gamma_{\widehat{\lambda}}(X_{\mathrm{test}})$ depends on both $X_{\mathrm{test}}$ and the entire calibration dataset.)
We now provide more formal details on the assumptions, the CRC method, and its statistical guarantee. Throughout the paper, the recurring assumptions on the model and data are the following.
Assumption 1.
The calibration and test data are observations of $n+1$ random variables $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X_{\mathrm{test}},Y_{\mathrm{test}})$ that are independent and identically distributed. Moreover, the predictor $f$ is deterministic.
Note that in practice, $f$ is typically the result of a training step, using another dataset $\mathcal{D}_{\mathrm{train}}$. In that case, the appropriate assumption is that the $n$ calibration points together with the test sample be independent and identically distributed conditionally on $\mathcal{D}_{\mathrm{train}}$. This implies that the CRC procedure could use calibration and test data from a different distribution than seen in training, so that one may reuse pre-trained models on a new specific task [14, 15].
The CRC method recalled next only handles a scalar parameter $\lambda\in\Lambda$, where $\Lambda\subset\mathbb{R}$. We also suppose that $\Lambda$ is compact (for simplicity), as well as the following.
Assumption 2.
The random loss functions $L_{i}:\Lambda\rightarrow\mathbb{R}$ defined by $L_{i}(\lambda)=\ell(Y_{i},\Gamma_{\lambda}(X_{i}))$ are non-increasing and right-continuous in $\lambda$, take their values in $\left[0,B\right]$ and satisfy $L_{i}(\bar{\lambda})=0$ almost surely, where $\bar{\lambda}=\max\Lambda$.
The monotonicity assumption on the $L_{i}$ is immediately satisfied when the prediction sets are nested (i.e. $\Gamma_{\lambda}(x)\subset\Gamma_{\lambda^{\prime}}(x)$ for all $\lambda\leq\lambda^{\prime}\in\Lambda$ and $x\in\mathcal{X}$) and when the loss function $\ell(y,S)$ is non-increasing in $S$ (i.e., $\ell(y,S)\geq\ell(y,S^{\prime})$ for all $S\subset S^{\prime}$). The earliest methods in Conformal Prediction correspond to the special case $\ell(y,S)=\mathbbm{1}_{y\notin S}$. Another example is the Recall of False Negative Rate loss. In Section IV we will use both examples for classification and localization tasks respectively, as well as other ones.
The objective is then to correctly tune the parameter $\lambda$ in order to control the behavior of $L_{\mathrm{test}}(\lambda)=\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$ with respect to a user-defined threshold $\alpha$. The CRC method of [53] computes
$\widehat{\lambda}=\inf\left\{\lambda\in\Lambda:\frac{1}{n+1}\sum\limits_{i=1}^{n}L_{i}(\lambda)+\frac{B}{n+1}\leq\alpha\right\}.$ (1)
Put differently, CRC finds the smallest $\lambda$ (i.e., the one leading to the smallest prediction set) such that the average loss $\frac{1}{n+1}\sum_{i=1}^{n}L_{i}(\lambda)+\frac{L_{\textrm{test}}(\lambda)}{n+1}$ over the $n+1$ calibration/test points is below the desired error rate $\alpha$, when anticipating a worst-case value of $B$ for the unseen test loss $L_{\textrm{test}}(\lambda)$.
Theorem 1 (Theorems 1 and 2 in [53]).
Under Assumptions 1 and 2, for any $\alpha\geq\frac{B}{n+1}$, the parameter $\hat{\lambda}$ defined in (1) satisfies
$\alpha-\frac{2B}{n+1}\leq\mathbb{E}\left[L_{\mathrm{test}}(\hat{\lambda})\right]\leq\alpha,$ (2)
where the lower bound holds under an additional assumption of continuity on the losses.
Importantly, the above expectation is over both $\mathcal{D}_{\text{cal}}$ and $(X_{\mathrm{test}},Y_{\mathrm{test}})$: roughly speaking, it is a mean computed over all possible joint observations of $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X_{\mathrm{test}},Y_{\mathrm{test}})$. There is no guarantee that (2) holds for all possible test instances, or for all calibration data sets. In particular, when $\ell(y,S)=\mathbbm{1}_{y\notin S}$, the event $\{Y_{\mathrm{test}}\notin\Gamma_{\lambda}(X_{\mathrm{test}})\}$ might occur because of an atypical test image $X_{\mathrm{test}}$, label $Y_{\mathrm{test}}$, or of atypical calibration points $(X_{i},Y_{i})$ in $\mathcal{D}_{\text{cal}}$. Ideally, one may hope for a guarantee that holds conditionally on each test instance, but such test-conditionality is impossible to obtain without further assumptions, as shown in [55].
Other Conformal Procedures. Multiple extensions exist to solve variations of this problem. For example, with the CLCP method of [56], one can replace the guarantee on the mean (expectation) with a guarantee on the median or quantiles of the loss. We may also want to obtain guarantees that hold not only on average over all classes, but on average per class [55, 57]. This can be useful to reduce over-correction or under-correction of some target classes. Furthermore, Conformal Prediction has been studied in other settings, such as Full CP [13], which trains the predictor and calibrates prediction sets on the same data, or the jackknife (leave-one-out) and cross-validation variants [58, 59, 60].
III-B Sequential Conformal Risk Control with 1+2 parameters
In the classical setting recalled in Section III-A, we control a loss on a prediction set indexed by a single parameter $\lambda$. Next, in order to match the object detection setting of Section II, we consider three parameters: $\lambda^{\mathrm{cnf}}$ for confidence thresholding, $\lambda^{\mathrm{loc}}$ for the localization task, and $\lambda^{\mathrm{cls}}$ for the classification task. The localization and classification tasks are dependent on the confidence task, and existing theory for conformal prediction cannot handle this situation. Therefore we introduce Sequential Conformal Risk Control (SeqCRC), a new method that extends the CRC algorithm to handle 1+2 sequentially dependent parameters, which is well suited for OD tasks.
Notation and assumptions. We denote by $L_{i}^{j}$, with $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$, the loss functions for the confidence ($L^{\mathrm{cnf}}$), localization ($L^{\mathrm{loc}}$) and classification ($L^{\mathrm{cls}}$) tasks. For each task $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$ we also write $\lambda^{j}\in\Lambda^{j}$ for the associated parameter and $\alpha^{j}$ for the desired error rate. We assume that the $\Lambda^{j}\subset\mathbb{R}$ are compact, and set $\bar{\lambda}^{j}=\max\Lambda^{j}$. (Our examples in Section IV are of the form $\Lambda^{j}=\bigl{[}0,\bar{\lambda}^{j}\bigr{]}$.)
Our procedure proceeds in two steps:
  • First is the confidence step, with the loss $L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$ depending only on a single parameter (as in Section III-A).
  • Then, two independent steps are carried out in parallel, those of localization and classification: they depend on the first one and are therefore denoted by $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$, where $\bullet\in\{\mathrm{loc},\mathrm{cls}\}$.
Though not written explicitly in this section, the random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are obtained through task-specific losses and predictions sets defined in Section IV, roughly of the form:888 Though not written explicitly below, the loss $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$ also depends on a matching between ground-truth and predicted boxes. See Section IV-A for details.
$L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})=\ell^{\mathrm{cnf}}\bigl{(}Y_{i},\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})\bigr{)}$
and
$L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})=\ell^{\bullet}\bigl{(}Y_{i},\Gamma^{\bullet}_{\lambda^{\mathrm{cnf}},\lambda^{\bullet}}(X_{i}))\;.$
For the analysis, we will assume the following two-step counterpart of Assumption 2. Recall that the $\Lambda^{j}\subset\mathbb{R}$ are compact, and that $\bar{\lambda}^{j}=\max\Lambda^{j}$.
Assumption 3.
The random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are non-increasing and right-continuous in all of their parameters,999i.e., non-increasing and right-continuous in each of the parameters, all the other parameters being fixed. (This should hold true for every sample path.) and take their values in $\left[0,B^{\mathrm{cnf}}\right]$ and $\left[0,B^{\bullet}\right]$ respectively. Also assume that for any $i\in\{1,\cdots,n\}$, we have $L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ almost surely.
In general, we do not assume that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})=0$ almost surely (this is not necessarily the case in Section IV), though this additional property implies additional guarantees, as shown in Theorem 2 below.
III-B1 The SeqCRC method
We now introduce our method; see Steps 1 and 2 below. We use the following notation for the empirical risks related to the confidence, localization, and classification tasks:
$R^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$ $=\frac{1}{n}\sum_{i=1}^{n}L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})\,,$
$R^{\bullet}_{n}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$ $=\frac{1}{n}\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})\,.$
We also use a (slightly) conservative variant of $R^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$ that anticipates the other two tasks:
$\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$ $=\max\Big{\{}R^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}}),R^{\mathrm{loc}}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}}),$ (3)
$\qquad\qquad R^{\mathrm{cls}}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})\big{\}}\,,$
where $\bar{\lambda}^{\mathrm{loc}}=\max\Lambda^{\text{loc}}$ and $\bar{\lambda}^{\mathrm{cls}}=\max\Lambda^{\text{cls}}$. Note that $\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})\in[0,\tilde{B}^{\mathrm{cnf}}]$ with $\tilde{B}^{\mathrm{cnf}}=\max\{B^{\mathrm{cnf}},B^{\mathrm{loc}},B^{\mathrm{cls}}\}$.
Step 1: For the confidence step, we compute:
$\lambda^{\mathrm{cnf}}_{+}$ $=\inf\left\{\lambda^{\mathrm{cnf}}\!\in\!\Lambda^{\mathrm{cnf}}\!:\!\frac{n\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})}{n+1}\!+\!\frac{\tilde{B}^{\mathrm{cnf}}}{n+1}\!\leq\!\alpha^{\mathrm{cnf}}\!\right\},$ (4)
$\lambda^{\mathrm{cnf}}_{-}$ $=\inf\left\{\lambda^{\mathrm{cnf}}\!\in\!\Lambda^{\mathrm{cnf}}\!:\!\frac{n\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})}{n+1}\!+\!\frac{0}{n+1}\!\leq\!\alpha^{\mathrm{cnf}}\!\right\},$ (5)
where we use the convention $\inf\varnothing=\max\Lambda^{\text{cnf}}=\bar{\lambda}^{\mathrm{cnf}}$.
Importantly, contrary to the CRC algorithm, we build two estimators instead of one: $\lambda^{\mathrm{cnf}}_{+}$ is used on test images for confidence thresholding, while $\lambda^{\mathrm{cnf}}_{-}$ is used in the second step (6) below. The reason we work with $\tilde{R}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}})$ instead of $R^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}})$ is to ensure that this second step is feasible, as will be clear from the analysis.
Step 2: For the second step (localization and classification), our SeqCRC method computes:
$\lambda^{\bullet}_{+}$ $=\inf\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{nR^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})}{n+1}+\frac{B^{\bullet}}{n+1}\leq\alpha^{\bullet}\right\}.$ (6)
III-B2 Theoretical analysis of SeqCRC
We now state our main theoretical result. It implies a probabilistic guarantee for the localization and classification tasks, and an additional guarantee for the confidence task under an additional assumption on the confidence loss function.
Theorem 2.
Suppose that Assumptions 1 and 3 hold true. [^10] To be more rigorous, in addition to Assumption 1, we should assume that $(L^{\text{cnf}}_{i},L^{\text{loc}}_{i},L^{\text{cls}}_{i})$ is fully determined (in a measurable way) by $(X_{i},Y_{i})$, as in the examples of Section IV. This implies the following useful property: that the random $3$-tuples $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$ are independent and identically distributed. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
$\mathbb{E}\left[L_{\mathrm{test}}^{\bullet}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\right]\leq\alpha^{\bullet}\;.$ (7)
Under the additional assumption that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i\in\{1,\ldots,n+1\}$,111111 Recall that $\bar{\lambda}^{\mathrm{cnf}}=\max\Lambda^{\mathrm{cnf}}$. we also have
$\mathbb{E}\left[L_{\mathrm{test}}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{+})\right]\leq\alpha^{\mathrm{cnf}}\;.$ (8)
Proof.
See Appendix B. ∎
The proof follows from similar arguments to the (single-step) CRC algorithm of [53], with some additional subtleties due to the sequential nature of the problem. In particular, the modified empirical risk $\tilde{R}^{\mathrm{cnf}}_{n}$ (3) in the first step is useful to ensure that the set in (6) in nonempty, and the “optimistic” estimator $\lambda^{\mathrm{cnf}}_{-}$ proves key for symmetry purposes.
The bound in (7) implies two individual guarantees, for the localization and classification tasks respectively. In practice, we usually want a joint guarantee on the two tasks: an example is given by the following immediate corollary (we use $\max\{a,b\}\leq a+b$ for $a,b\geq 0$).
Corollary 1.
Under the same assumptions as in Theorem 2,
$\mathbb{E}\left[\max\big{(}L^{\mathrm{loc}}_{\mathrm{test}}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+}),L^{\mathrm{cls}}_{\mathrm{test}}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cls}}_{+})\big{)}\right]\leq\alpha^{\mathrm{tot}},$ (9)
where $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{loc}}+\alpha^{\mathrm{cls}}$.
In other words, it is possible to get a global guarantee at a level $\alpha^{\mathrm{tot}}$ by splitting it between the localization and classification tasks. The same logic is applicable with three terms, if need be to include the confidence loss in the final guarantee (at the cost of a larger total error $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{cnf}}+\alpha^{\mathrm{loc}}+\alpha^{\mathrm{cls}}$). For instance, simultaneously guaranteeing a sufficient number of predictions while ensuring their localization and classification. Importantly, the choices of $\alpha^{\mathrm{loc}}$ and $\alpha^{\mathrm{cls}}$ given a global $\alpha^{\mathrm{tot}}$ can be decided according to the difficulty and practical importance of each task, as long as they are not done in a data-dependent way; that is, one cannot choose the values of $\alpha$ after observing the calibration data.
We now comment on the assumptions from a practical standpoint.
On the vanishing loss assumption. In Assumption 3 the two conditions $L^{\mathrm{loc}}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})=0$ and $L^{\mathrm{cls}}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})=0$ on the localization and classification losses have practical implications. When applied to our examples in Section IV, they correspond to assuming that, after Non-Maximum Suppression, there is at least one predicted object per image if the underlying matching is non-injective (i.e., multiple ground-truth objects can be matched to one predicted object), or at least as many predictions as ground truths if it is injective (e.g., via Hungarian matching, one can match every ground-truth object to its own prediction).
On the monotone loss assumption. Though intuitive, the assumption that the loss functions are monotone can fail in practice for $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$. While it may not appear obvious from the loss definitions in the next section, the matching required for their computation is the root of this issue (see Remarks 1 and 2 for details). This implies that the localization and classification loss functions are typically non-monotone in the $\lambda^{\mathrm{cnf}}$ parameter, even if they are close to monotone in most of our experiments.
To avoid this (small) gap between theory and practice, in Section IV-E, we propose a simple monotonization trick: within (4), (5), and (6) we replace $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$ with tight upper bounds that are provably monotone, as done similarly in [53]. These modified loss functions can be computed on the fly, and preserve the guarantees (7) and (8). Nevertheless, this comes at a cost, as these conservative losses lead to larger parameters $\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ and therefore larger prediction sets, as the gap to monotonicity of the loss increases. This gap should therefore be as small and localized as possible, especially for values below $\alpha$.
IV Application-tailored Prediction Sets & Losses
The general method presented in the previous section can be instantiated with various losses and prediction sets to meet specific practical requirements. This section presents a selection of such sets and losses and compares their features. They are introduced for confidence, localization and classification tasks respectively in Sections IV-B, IV-C and IV-D. Additional losses and prediction sets are discussed in Appendix E. Finally, in Section IV-E we present algorithms with their high-level pseudo-codes both for calibration (where we monotonize irregular losses on the fly, in order to meet the assumptions behind SeqCRC) and for inference.
Throughout this section, we work with losses that are $[0,1]$-valued, i.e., $B^{\mathrm{cnf}}=1$ and $B^{\bullet}=1$. Moreover we define prediction sets and losses for a deterministic example consisting in an image $x$ and its ground truth $y$. (We thus drop the index $i$ in $L_{i}$, except when multiple images are considered, e.g., in the algorithms’ pseudo-code.)
Next, we start by discussing essential preliminary considerations on two topics that are specific to object detection: the level of the statistical guarantee and matching.
IV-A Level of Guarantee & Matching
Although object detection might naively be interpreted as a combination of two simple problems—classification and regression—there are modeling and algorithmic subtleties that are key to make SeqCRC’s theoretical guarantee (Theorem 2) meaningful in OD practice.
First, the expectations in (7) and (8) are averages over “instances”, whose definition is ambiguous in the OD context. We find two alternatives:
  1. 1.
    Object-level guarantee (instances = objects): Each distinct annotated object in an image is considered an instance. For a dataset of $n$ images, where image $X_{i}$ contains $\lvert Y_{i}\rvert$ objects, the total number of instances is $\sum_{i=1}^{n}\lvert Y_{i}\rvert$.
  2. 2.
    Image-level guarantee (instances = images): Each image and its associated set of objects in the dataset is considered a single instance. For a dataset of $n$ images, this results in $n$ instances.
The two settings are strongly connected, and while the former is closer to the usual approach in OD, we focus on the more general image-level one in the core of our article. We argue that the image-level setting allows for more flexibility in the design of losses, and in particular to account for interactions between objects on an image (more details in Appendix A).
Second, in order to compute losses on calibration data and metrics on test data, we need to define a rule for matching the predictions of the model $f$ to ground-truth annotations: in our setting, an image can have multiple objects (boxes) and in practice, it is not trivial to determine which predicted box is associated with which ground-truth box, if any at all. This matching (ground-truth $\rightarrow$ prediction) does not need to be injective (but may be, such as using a Hungarian algorithm), meaning that multiple ground-truth boxes can be matched to the same predicted box.
Hereon, we assume that every image $x$ produces a matching $\pi_{x}$ that associates each ground-truth $(b_{j},c_{j})$ box/class pair with its nearest prediction $(b_{\pi_{x}(j)},c_{\pi_{x}(j)})$ with respect to a “distance” function $d$ [^12] We use the term “distance” even if $d$ is not necessarily a metric from a mathematical viewpoint.
The choice of $d$ is critical, as it shapes the prediction sets obtained using our conformal procedure. A satisfactory distance, which accurately connects true and predicted objects, is hard to design: there is a trade-off between minimizing spatial distance, and distance in the classification space. Indeed, for a given ground truth, the closest prediction spatially can be very incorrect in classification, and the closest prediction in classification can be far spatially. For localization, following the seminal work of [14], we use the asymmetric signed Hausdorff distance (see also [61]):
$d_{\mathrm{haus}}(b,\hat{b})=\max\left\{\hat{b}^{\leftarrow}-b^{\leftarrow},\hat{b}^{\uparrow}-b^{\uparrow},b^{\rightarrow}-\hat{b}^{\rightarrow},b^{\downarrow}-\hat{b}^{\downarrow}\right\}.$
Intuitively, it corresponds to the smallest margin (in pixels) to add to any of the predicted coordinates of $\hat{b}$ such that the extended box covers the ground-truth box entirely, aligning with localization objectives.
For classification, we propose using the nonconformity score of [57] as a distance, which we refer to as LAC, from their Least Ambiguous Classifier:
$d_{\mathrm{LAC}}(c,\hat{c})=1-\hat{c}_{c},$
where $c$ is the integer true-class label and $\hat{c}$ a probability vector (e.g. softmax output). It measures how far the predicted probability $\hat{c}$ for the true class $c$ is from being $1$. We therefore propose to merge the two, balanced by a parameter $\tau\in(0,1)$, resulting in:
$d_{\mathrm{mix}}\left((b,c),(\hat{b},\hat{c})\right)=\tau d_{\mathrm{LAC}}(c,\hat{c})+(1-\tau)d_{\mathrm{haus}}(b,\hat{b}).$ (10)
The distance functions $d_{\mathrm{LAC}}$ and $d_{\mathrm{haus}}$ are natural candidates since they capture identical or similar phenomena compared to the localization and classification losses used in the next sections. Other distance choices are possible, but not all of them are relevant. For instance, we also consider the Generalized IoU (Intersection over Union) distance function, following a common practice in OD [62]:
$d_{\mathrm{GIoU}}(b,\hat{b})=1-\frac{\mathrm{area}(b\cap\hat{b})}{\mathrm{area}(b\cup\hat{b})}+\frac{\mathrm{area}(\tilde{b}\backslash(b\cup\hat{b}))}{\mathrm{area}(\tilde{b})},$
where $\tilde{b}$ is the smallest convex hull enclosing both $b$ and $\hat{b}$. As we will see in Section V-C, this distance fails in practice compared to the other propositions.
We now address the confidence, localization and classification tasks. Recall that throughout this section we work with a generic example $(x,y)$, which can belong to calibration or test data. The notation $x$ stands for the input image, while $y$ denotes the list (or sequence) of all annotated ground-truth objects on $x$, which are elements of ×$\mathcal{B}\times\mathcal{C}$ (bounding box and class, see Section II-A).
IV-B First Step: Confidence
For the first step introduced in Sec. III-B1, and as (2) in Fig. 1 we define our prediction set and loss as follows.
Prediction set
For an input image $x$, the output of our OD model $f$ followed by NMS is a set of $N^{\mathrm{nms}}$ objects (see Sec. II-A and Fig. 2) with confidence scores $o(x)_{1},\dots,o(x)_{N^{\mathrm{nms}}}$ in $[0,1]$. In the object detection literature, trained models are evaluated using mean Average Precision (mAP), which is computed by generating precision-recall curves for each class with a range of thresholds on the confidence scores. For deployment, users must choose a fixed threshold (e.g., 0.5); this threshold determines how many predictions are excluded from the final set of detections. We offer there to calibrate this threshold, in such a way that we obtain a statistical guarantee on a specific notion of error important for the user.
More formally, let $\lambda^{\mathrm{cnf}}\in\Lambda^{\mathrm{cnf}}=[0,1]$. We start by identifying objects that satisfy
=$\mathcal{I}=\bigl{\{}k\in\{1,\cdots,N^{\mathrm{nms}}\}:o(x)_{[k]}\geq 1-\lambda^{\mathrm{cnf}}\bigr{\}}\;,$
where $o(x)_{[k]}$ is the $k$-th largest among the $\{o(x)_{k}\}_{k}$. Then, for any image $x$, the (post-processed) prediction set $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ consists of all objects with sufficiently high confidence scores:
$\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)=\bigl{\{}(\hat{b}(x)_{[k]},\hat{c}(x)_{[k]})\bigr{\}}_{k\in\mathcal{I}}\;.$ (11)
Losses
Since the loss $L^{\mathrm{cnf}}$, which controls the amount of bounding boxes, comes before the localization step, which controls their size, it follows that $L^{\mathrm{cnf}}$ controls a trade-off between the amount and size of bounding boxes through the definition of $L^{\mathrm{cnf}}$ and the choice of $\alpha^{\mathrm{cnf}}$. This choice should be made depending on operational and downstream objectives or requirements. We provide several examples.
A first natural approach could be to force the number $\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)\rvert$ of predictions to be close to the number $\lvert y\rvert$ of true boxes. However, this is not feasible because the loss would not be non-increasing in $\lambda^{\mathrm{cnf}}$. Therefore, we introduce the “box-count-threshold” loss, that controls this quantity in a one-sided manner, such that at least as many predictions as the number of true objects are kept:
$L^{\mathrm{cnf}}_{\mathrm{box-count-threshold}}(\lambda^{\mathrm{cnf}})=\begin{cases}0&\text{ if }\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)\rvert\geq\lvert y\rvert,\\ 1&\text{ otherwise}.\end{cases}$
This loss imposes a strong penalty because having only one less prediction than the number of ground truths yields a loss of $1$. The “box-count-recall” loss is a relaxed version:
$L^{\mathrm{cnf}}_{\mathrm{box-count-recall}}(\lambda^{\mathrm{cnf}})=\frac{\left(\lvert y\rvert-\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)\rvert\right)_{+}}{\lvert y\rvert},$
where $(\cdot)_{+}=\max(0,\cdot)$.
These losses do not account for the locations of the predictions and ground truths. We introduce and discuss additional examples of losses in Appendix E.
IV-C Second Step: Localization
Figure 3: Conformal Localization Step. The localization step, one of the two elements of (3) in Fig. 1, consists in applying corrective margins to each of the predicted bounding boxes.
For the second step introduced in Sec. III-B1 and illustrated in Fig. 3, recall that by “localization” we refer to how well a bounding box contains the ground-truth bounding box of an object, regardless of its class. Similarly to the previous confidence step, we work with sequences of bounding boxes (multiple objects per image).
Prediction sets
Roughly speaking, we alter the boxes obtained after the confidence step by modifying their coordinates. For an image $x$, the “prediction set” $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)$ is a sequence of boxes with the same size as $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$; it contains the same bounding boxes as $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ but adjusted as in Eqs. (12) or (13) below.
More precisely, let $\hat{b}_{k}=(\hat{b}^{\leftarrow},\hat{b}^{\uparrow},\hat{b}^{\rightarrow},\hat{b}^{\downarrow})$ denote the $k$-th box in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$. (We ignore the second element of the tuple in (11), which is treated in Section IV-D below.) Following [14, 16], a natural way to adjust the predicted box $\hat{b}_{k}$ is to add an “additive” margin (in pixels) around it. This new box is the $k$-th element $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)_{k}$ of our prediction set $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)$:
$\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)_{k}=\hat{b}_{k}+\begin{pmatrix}-\lambda^{\mathrm{loc}}&-\lambda^{\mathrm{loc}}&\lambda^{\mathrm{loc}}&\lambda^{\mathrm{loc}}\end{pmatrix},$ (12)
where $\lambda^{\mathrm{loc}}\in\Lambda^{\mathrm{loc}}\subset\mathbb{R}_{+}$, usually bounded by the size of the image. However, all bounding boxes will be expanded by an equal margin, which can be disproportional to their size: predictions for small objects may suffer very large corrections and vice versa. To mitigate this issue, one can instead build a “multiplicative” margin as in [14, 16]: one multiplies the parameter $\lambda^{\mathrm{loc}}$ by factors $\hat{w}$ and $\hat{h}$, respectively corresponding to the height and width of the predicted box:
$\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)_{k}=\hat{b}_{k}+\lambda^{\mathrm{loc}}\begin{pmatrix}-\hat{w}_{k}&-\hat{h}_{k}&\hat{w}_{k}&\hat{h}_{k}\end{pmatrix},$ (13)
where $\lambda^{\mathrm{loc}}\in\Lambda^{\mathrm{loc}}\subset\mathbb{R}_{+}$, also bounded by the size of the image.
Note that, as already done for regression tasks (in, e.g., [50, 63, 54]), we could replace these width and height factors by any notion of local uncertainty to improve the adaptivity of our conformal correction (see Appendix E). It is also possible to obtain corrections that depend on the class associated with the box [23].
Losses
For a given image $x$ with ground-truth labeling $y$, let $b_{j}$ denote the $j$-th bounding box in $y$. Following Section IV-A, in order to define the losses, we compute a matching $\pi_{x}$ that maps indices $j$ of true bounding boxes $b_{j}$ to indices $k$ of predicted boxes $\hat{b}_{k}$ in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$. The predicted box paired to a true box $b_{j}$ is thus denoted by $\hat{b}_{\pi_{x}(j)}$.
A natural loss introduced by [14] consists in checking whether a proportion at least $\tau\in[0,1]$ of true boxes $b_{j}$ on an image $x$ are entirely covered by their matched predictions $\hat{b}_{\pi_{x}(j)}$ after resizing the latter as in (12) or (13). Writing $\hat{b}_{k}^{\lambda^{\mathrm{loc}}}=\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)_{k}$ for simplicity, this reads:
$L^{\mathrm{loc}}_{\mathrm{thr}}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}})=\begin{cases}0\text{ if }\dfrac{\lvert\{b_{j}\in y,b_{j}\subseteq\hat{b}_{\pi_{x}(j)}^{\lambda^{\mathrm{loc}}}\}\rvert}{\lvert y\rvert}\geq\tau\\[8.61108pt] 1\text{ otherwise}.\end{cases}$
This has two drawbacks. First, the choice of $\tau$ must be done before looking at the data, and may significantly impact the result (a small change of $\tau$ may cause a large variation in the loss). Second, it penalizes without distinction boxes that fail to be covered by a single pixel, or by a significant margin.
For this purpose, following [16, 15], we also study two relaxations of the previous loss. The first consists in directly controlling the previous proportion, and is therefore referred to by box-wise recall:
$L^{\mathrm{loc}}_{\rm box}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}})=1-\frac{\lvert\{b_{j}\in y,b_{j}\subseteq\hat{b}_{\pi_{x}(j)}^{\lambda^{\mathrm{loc}}}\}\rvert}{\lvert y\rvert}\;.$ (14)
However, the inclusion condition remains restrictive. We can further relax the loss using the average proportion of the covered areas of the true boxes $b_{j}$, via the pixel-wise loss:
$L^{\mathrm{loc}}_{\rm pix}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}})=1-\frac{1}{\lvert y\rvert}\sum_{b_{j}\in y}\frac{\mathrm{area}(b_{j}\cap\hat{b}_{\pi_{x}(j)}^{\lambda^{\mathrm{loc}}})}{\mathrm{area}(b_{j})}\;,$
where $\mathrm{area}(b)=(b^{\rightarrow}-b^{\leftarrow})(b^{\downarrow}-b^{\uparrow})$ is the area of a bounding box, and $b_{j}\cap\hat{b}$ is the (possibly degenerate) bounding box obtained by intersecting $b_{j}$ and $\hat{b}$.
Remark 1 (On the monotonicity assumption).
In Assumption 3 we require that losses be non-increasing in the parameters $\lambda^{\mathrm{loc}}$ and $\lambda^{\mathrm{cnf}}$. While the three examples $L^{\mathrm{loc}}_{\mathrm{thr}},L^{\mathrm{loc}}_{\rm box},L^{\mathrm{loc}}_{\rm pix}$ are easily seen to be non-increasing with respect to the additive/multiplicative margin $\lambda^{\mathrm{loc}}$, the dependency in $\lambda^{\mathrm{cnf}}$ is not necessarily monotone.
Figure 4: Monotonization of a Localization Loss. The plot represents the box-wise recall loss (Eq. (14)) as a function of the confidence parameter $\lambda^{\mathrm{cnf}}$. The various curves (colors) correspond to different fixed values of $\lambda^{\mathrm{loc}}$. The solid lines are the raw losses, and the dotted ones are their monotonized variants. While we could expect raw losses to be non-increasing, they are usually not, due to (partially) diverging goals between the localization loss and the distance $d$ used in the matching algorithm.
See Fig. 4 for an illustration. The reason is that, as $\lambda^{\mathrm{cnf}}$ increases, more and more boxes are added in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$, so that the matching $\pi_{x}$ between true boxes $b_{j}$ and boxes $\hat{b}_{k}$ in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ changes. This change is guided by the “distance” function $d$ used to compute $\pi_{x}$, which is only vaguely related to the three loss examples above. See Appendix E-D for additional details.
In Section IV-E we will introduce a monotonization trick to fix this issue.
IV-D Classification
We now address the classification task, as illustrated in Fig. 5. We adapt two well-known CP methods for classification to our setting: the Least Ambiguous Set-Valued Classifiers (LAC) of [57] and that of [64], which we refer to as Adaptive Prediction Sets (APS). However, other methods from the literature could be applied [65, 66, 67, 68].
Prediction sets
Our prediction sets for the classification task are also defined element-wise (for every index $k$). We let $\hat{c}_{k}$ denote the $k$-th classification prediction within $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$, ignoring the previously handled location component. Recall from Section II-A that $\hat{c}_{k}$ is a probability vector (e.g., a softmax output) over the label space $\mathcal{C}=\{1,\cdots,K\}$.
Letting $\lambda^{\mathrm{cls}}\in\Lambda^{\mathrm{cls}}=[0,1]$, the LAC prediction set [57] is defined by
$\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{k}=\left\{\kappa\in\mathcal{C}:\hat{c}_{k}(\kappa)\geq 1-\lambda^{\mathrm{cls}}\right\},$ (15)
where $\hat{c}_{k}(\kappa)$ is the softmax score of the $k$-th prediction for the class $\kappa$. Put differently, all sufficiently likely classes (according to the softmax prediction) are predicted.
Figure 5: Conformal Classification Step. The classification step, one of the two elements of (3) in Fig. 1, consists in a predicting a set of labels for each object, with the guarantee to contain the ground truth with high probability.
The APS variant by [64] consists in retaining the most likely classes too, but in a way that the cumulative softmax probability is above a given threshold. More precisely, one first ranks the softmax scores in decreasing order: let $\kappa_{[1]},\cdots,\kappa_{[K]}\in\mathcal{C}$ be such that $\hat{c}_{k}(\kappa_{[1]})\geq\cdots\geq\hat{c}_{k}(\kappa_{[K]})$. (The order can be different for different values of $k$.) The APS prediction set is then defined by [^13] When $\lambda^{\mathrm{cls}}=1$, we set $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{k}=\mathcal{C}$ by convention.
$\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{k}=\left\{\kappa_{[1]},\dots,\kappa_{[\hat{m}(\lambda^{\mathrm{cls}})]}\right\},$ (16)
where
$\hat{m}(\lambda^{\mathrm{cls}})=\min\left\{m:\sum_{l=1}^{m}\hat{c}_{k}(\kappa_{[l]})>\lambda^{\mathrm{cls}}\right\}.$
Losses
Let $c_{j}\in\mathcal{C}$ denote the class label of the $j$-th object in the ground truth $y$. We evaluate a binary loss for each true object, by checking whether its class $c_{j}$ belongs to the associated prediction set component $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{\pi_{x}(j)}$. We then aggregate these loss values. In our experiments we compute an average over all true objects, but some alternatives are discussed in Appendix A. The average version reads:
$L^{\mathrm{cls}}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}})=\frac{1}{\lvert y\rvert}\sum_{c_{j}\in y}\mathbbm{1}_{c_{j}\notin\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{\pi_{x}(j)}}\;.$ (17)
Remark 2.
Similarly to Remark 1, for the two prediction sets defined in (15) and (16), the loss $L^{\mathrm{cls}}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}})$ is non-increasing in $\lambda^{\mathrm{cls}}$ but not necessarily in $\lambda^{\mathrm{cnf}}$.
IV-E Algorithm
In this section we introduce—in a general form—the algorithms used to compute the prediction sets and tune their parameters. They can be used for any choices of prediction set and losses compatible with our method.
Algorithm 1 SeqCRC: Calibration of Parameters
1:Calibration dataset $\mathcal{D}_{\text{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, object detector $f$ (including NMS), matching procedure $\pi$, desired error rates $\alpha^{\mathrm{cnf}},\alpha^{\mathrm{loc}},\alpha^{\mathrm{cls}}$.
2:Calibrated parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cnf}}_{-},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$
3:Step 0: Definitions
4:Define prediction sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}},\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}},\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$
5:Define losses $L_{i}^{\mathrm{cnf}},L_{i}^{\mathrm{loc}},L_{i}^{\mathrm{cls}}$ for confidence, localization, and classification
6:Step 1: Compute predictions
7:for $i=1$ to $n$ do
8:     $\hat{Y}_{i}\leftarrow f(X_{i})$
9:end for
10:Step 2: Calibrate confidence parameters
11:$\lambda^{\mathrm{cnf}}_{+}\leftarrow$ Solve Eq. (4) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, the sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$, $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ , and the losses $L_{i}^{\mathrm{cnf}}$, $L_{i}^{\mathrm{loc}}$, $L_{i}^{\mathrm{cls}}$ using Algorithm 3 (with the monotonization trick)
12:$\lambda^{\mathrm{cnf}}_{-}\leftarrow$ Solve Eq. (5) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, the sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$, $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ , and the losses $L_{i}^{\mathrm{cnf}}$, $L_{i}^{\mathrm{loc}}$, $L_{i}^{\mathrm{cls}}$ using Algorithm 3 (with the monotonization trick)
13:Step 3: Compute matchings (see Section IV-A)
14:for $i=1$ to $n$ do
15:     $\hat{Y}_{i}\leftarrow f(X_{i})$
16:     $\pi_{x}\leftarrow\pi(Y_{i},\hat{Y}_{i})$
17:end for
18:Step 4: Calibrate localization parameter
19:$\lambda^{\mathrm{loc}}_{+}\leftarrow$ Solve Eq. (6) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, $L_{i}^{\mathrm{loc}}$ and $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ using Algorithm 4 (with the monotonization trick)
20:Step 5: Calibrate classification parameter
21:$\lambda^{\mathrm{cls}}_{+}\leftarrow$ Solve Eq. (6) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, $L_{i}^{\mathrm{cls}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ using Algorithm 4 (with the monotonization trick)
In Algorithm 1 we present a high-level pseudo-code to compute the parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cnf}}_{-},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ defined in (4), (5) and (6), while meeting the monotonicity requirements on the loss functions $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$. (Recall from Section IV, Remarks 1 and 2, that these losses are not necessarily monotone in $\lambda^{\mathrm{cnf}}$.) As mentioned in Section III-B, this is achieved via a monotonization trick (e.g., [53]), i.e., by replacing $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$ with tight upper bounds that are provably monotone. More specifically, Algorithm 1 uses two monotonization subroutines: Algorithm 3 for the confidence step, and Algorithm 4 for the localization/classification step, presented in Appendix F.
These two algorithms “monotonize” the loss on the fly, to improve computational and memory efficiency. Due to their complexity and their implementation not being straightforward, more details are given in Appendix F.
Requirements In these algorithms, we assume the user has already trained an object detector $f$. We further suppose that they have chosen losses, prediction sets, error levels for all three steps, and a matching procedure $\pi$, as described in Sections IV-A through IV-D.
Edge cases We must also handle two particular edge cases: an image may contain no ground-truth objects (i.e., $y$ is empty), or no predictions after the confidence thresholding as in Eq. (11). In the first case, we set the loss to zero, and in the second case, to $B$ (in this section, for all losses we have $B=1$).
Algorithm 2 contains the procedure to follow for an inference on a test image. Using the same prediction set definitions as in calibration, and the parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ obtained from it, it builds prediction sets that satisfy the conformal guarantee of Theorem 2. The process is considerably simpler than calibration, as it solely involves constructing the prediction sets. This procedure is computationally inexpensive for all the prediction sets and losses introduced and can be regarded as negligible when compared to the cost of a forward pass of $f$. Importantly, only the $\lambda_{+}$ values are used during inference.
Algorithm 2 SeqCRC: Inference
1:Test image $X_{\mathrm{test}}$, object detector $f$ (including NMS), calibrated parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ and defined prediction sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}},\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}},\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$
2:Prediction sets $\Gamma^{\mathrm{cnf}},\Gamma^{\mathrm{loc}},\Gamma^{\mathrm{cls}}$
3:Step 1: Confidence filtering
4:$\hat{Y}_{\mathrm{test}}\leftarrow f(X_{\mathrm{test}})$
[…]
6:Step 2: Localization and classification prediction sets
7:for $k$ in $1\dots\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}_{+}}(X_{\mathrm{test}})\rvert$ do
8:     Compute $\Gamma^{\mathrm{loc}}\leftarrow\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+}}(X_{\mathrm{test}})_{k}$ as chosen in Section IV-C
9:     Compute $\Gamma^{\mathrm{cls}}\leftarrow\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cls}}_{+}}(X_{\mathrm{test}})_{k}$ as chosen in Section IV-D
10:end for
V Experiments
In this section, we present a comprehensive set of experiments designed to assess the performances of our SeqCRC approach to Object Detection (OD). These experiments include large-scale evaluations that result in a thorough benchmark, encompassing all matchings, losses, and prediction sets introduced in Section IV-A through IV-D.
Our goal is to provide insight into the practical interpretation and implications of our conformal method, highlighting both its strengths and limitations, while clarifying the practical process leading to statistical guarantees.
TABLE I: Comparison of set sizes and risks for all losses and prediction sets from Section IV, using DETR-50 as a reference model. The results are split into four categories: Confidence (Cnf.) and Localization (Loc.) losses, as well as Localization (Loc.) and Classification (Cls.) Prediction Sets (Pred. Sets). The risk levels are $\alpha^{\mathrm{cnf}}=0.02$ for the confidence task, $\alpha^{\mathrm{loc}}=0.05$ for localization, and $\alpha^{\mathrm{cls}}=0.05$ for classification. Thus, we apply the SeqCRC method at global level $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{loc}}+\alpha^{\mathrm{cls}}=0.1$ (see Corollary 1). For each setting, the results are shown using optimal choices for all remaining parameters.
Setting Set Size Task Risk Global Risk
Loss Cnf. box_count_threshold 26.379 0.019 0.084
box_count_recall 18.868 0.018 0.085
Loc. thresholded 1.626 0.045 0.097
boxwise 1.578 0.049 0.098
pixelwise 1.056 0.051 0.100
Pred. Sets Loc. additive 1.094 0.051 0.099
multiplicative 1.056 0.051 0.100
Cls. aps 1.072 0.051 0.083
lac 1.044 0.051 0.083
Figure 6: Histogram of set sizes for all combinations of losses, prediction sets and matching functions for $\alpha^{\mathrm{tot}}=0.1$, using the detector DETR. In blue the conformal set sizes in Localization, and in red, the classification ones (see Section V-A2). While the two have different units, their scale is similar, and, in distribution, both display a heavy tail. A large share of experiment settings lead to dramatically worse performance (= set size) at inference.
Figure 7: Example of a successful conformal prediction. The model’s predictions are close to the ground truths, and the conformal predictions for the localization and classification tasks both include their respective ground truths.
V-A Experimental Setup
In these experiments, we use two reference models: YOLOv8 [40] (in particular YOLOv8x [^14] https://github.com/ultralytics/ultralytics), a recent improvement in a long line of single-stage detectors including [38, 39], and DETR [43] (precisely, DETR-50151515https://github.com/facebookresearch/detr), a well-known transformer-based predictor. It is important to note that while all results here are based on these models, the method we present is model-agnostic and can be applied straightforwardly to virtually any other model (following our setting in Section II). The choice of these models emphasizes the broad applicability of our new framework, as well as the potential advantages or shortcomings of some families of models.
We employ the MS-COCO dataset [69], a standard benchmark in the field. For our calibration and inference procedures, we use the validation split of this dataset. This approach is common when an annotated dataset that is independent of training and model selection is unavailable. While potentially breaking Assumption 1, the practical availability constraints often make the validation split the best choice for conformal calibration and evaluation. The validation data is then split into two equal segments of $n=2500$ instances each, dedicated to calibration and inference respectively.
V-A1 Methods
We conduct experiments on MS-COCO for both models (YOLO8 and DETR) and for all combinations of losses, distance functions (matchings), and prediction sets introduced in Section IV. For the matching “mix” referring to distance (10), we set $\tau=0.25$, providing a fair balance between localization and classification distances, according to preliminary experiments. For all experiments, the IoU threshold for the NMS is set to $0.5$. We focus on the setting of Corollary 1 that offers a joint guarantee on the localization and classification tasks (hence $\alpha^{\mathrm{tot}}=\alpha^{\mathrm{cls}}+\alpha^{\mathrm{loc}}$ here). We consider two global error levels : $\alpha^{\mathrm{tot}}=0.1$ and $\alpha^{\mathrm{tot}}=0.2$. For the former we set $\alpha^{\mathrm{cnf}}=0.02$, $\alpha^{\mathrm{loc}}=0.05$ and $\alpha^{\mathrm{cls}}=0.05$ for confidence, localization and classification respectively. Similarly for the latter we set $\alpha^{\mathrm{cnf}}=0.03$, $\alpha^{\mathrm{loc}}=0.1$, $\alpha^{\mathrm{cls}}=0.1$. Finally, to mitigate the pathological second-step loss values towards $\lambda^{\mathrm{cnf}}=1$ (i.e. a null confidence threshold $1-\lambda^{\mathrm{cnf}}$) observed in Fig. 4, we filter out predictions with confidence lower than $10^{-3}$. This threshold is chosen independently of the data, for all experiments. (This thus corresponds to refining Steps 1 and 2 in Fig. 1, and does not break the guarantees of Theorem 2.) This leads to much lighter correction when applying the monotonization trick and therefore to smaller set sizes.
V-A2 Metrics
For evaluation, we follow a common practice in Conformal Prediction, that of reporting the two metrics of Empirical Risk and Set Size. We examine their relationships with OD metrics in Appendix C-B. Throughout this section, we focus on the test data sequence $(X_{\mathrm{test},i},Y_{\mathrm{test},i})_{i=1}^{n_{\mathrm{test}}}$.
Firstly, the empirical risk (or simply risk thereafter) is the empirical estimation of the validity of the procedure, that is, the average loss measured empirically on test data. (In the case of a binary loss, it corresponds to $1$ minus the coverage, one of the main metrics in Conformal Prediction.) For each task $j\in\{\mathrm{cnf,loc,cls}\}$, the empirical risk that we measure is
$j\text{-Risk:}\quad\left\{\begin{array}[]{ll}\frac{1}{n_{\mathrm{test}}}\sum\limits_{i=1}^{n_{\mathrm{test}}}L^{j}_{\mathrm{test,i}}(\lambda^{\mathrm{cnf}}_{+})&\mathrm{if~{}}j=\mathrm{cnf}\\ \frac{1}{n_{\mathrm{test}}}\sum\limits_{i=1}^{n_{\mathrm{test}}}L^{j}_{\mathrm{test,i}}(\lambda^{\mathrm{cnf}}_{+},\lambda^{j}_{+})&\mathrm{if~{}}j\in\{\mathrm{loc,cls}\}\end{array}\right.$
The $j$-Risk can be compared against the target value $\alpha^{j}$, for each task $j\in\{\mathrm{cnf,loc,cls}\}$ individually. However, a Global Risk can be computed by
$\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\max\left\{L^{\mathrm{loc}}_{\mathrm{test},i}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+}),L^{\mathrm{cls}}_{\mathrm{test},i}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cls}}_{+})\right\}.$
This corresponds to the joint risk controlled in Corollary 1 (therefore compared against $\alpha^{\mathrm{tot}}$), and it is our main target. As mentioned after the corollary, one may include the confidence risk if desired; in our experiments, we focus on localization and classification only.
Secondly, for each task (including confidence), we measure the mean set size, averaged over images. As the sets take various shapes depending on the tasks, we introduce three different metrics, tailored for interpretability.
First, set sizes of confidence sets are defined by
[…]
This metric measures the average number of boxes output by our procedure per image. Then, for localization, the metric we introduce is inspired from [15]; it is given by the ratio of the conformal box’ area over that of its original prediction (referred to as Stretch in the source). More precisely:
[…]
where $n_{\mathrm{test},i}=\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}_{+}}(X_{\mathrm{test},i})\rvert$ is the (final) number of predicted boxes on the $i$-th test image, and where $\hat{b}_{k}$ and $\hat{b}_{k}^{\lambda^{\mathrm{loc}}_{+}}$ are the $k$-th elements of $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}_{+}}(X_{\mathrm{test},i})$ and $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{loc}}_{+}}(X_{\mathrm{test},i})$ respectively—in other words, the same $k$-th predicted box before and after altering its size with the $\lambda^{\mathrm{loc}}_{+}$ parameter (e.g. an additive margin).
Finally, the set size for classification prediction sets is defined (similarly as above) by averaging over two levels: within a test image, and then over the test set;
[…]
where $\hat{c}_{k}^{\lambda^{\mathrm{cls}}_{+}}=\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cls}}_{+}}(X_{\mathrm{test},i})_{k}$ denotes the classification prediction set for the $k$-th prediction on image $X_{\mathrm{test},i}$.
TABLE II: Effect of target $\alpha^{\mathrm{tot}}$ level on set sizes for Confidence (Cnf.), Localization (Loc.) and Classification (Cls.) for four matching functions: the Generalized IoU (GIoU), the asymmetric signed Hausdorff divergence (Hausdorff), the Least-Ambiguous Classifier non-conformity score (LAC), and a mixture of the last two (Mix), the best performing overall.
Matching Target $\alpha^{\mathrm{tot}}$ Cnf. Size Loc. Size Cls. Size
GIoU 0.1 26.379 29.148 46.757
0.2 14.406 24.455 34.806
Hausdorff 0.1 26.379 1.056 44.271
0.2 14.406 1.000 23.466
LAC 0.1 26.379 15.239 1.044
0.2 23.974 8.314 0.669
Mix 0.1 26.379 1.396 8.709
0.2 23.974 1.043 0.906
TABLE III: Comparison of set sizes for all losses and prediction sets from Section IV for two models: DETR and YOLOX. The comparison is done for a global error level $\alpha^{\mathrm{tot}}=0.2$. The set size reported are the ones of the associated task in its part of the table.
Metric DETR YOLOv8
Risk Set Size Risk Set Size
Loss Cnf. box_count_threshold 0.027 23.974 0.022 15.412
box_count_recall 0.026 14.406 0.031 10.138
Loc. thresholded 0.096 1.444 0.090 4.779
boxwise 0.105 1.322 0.096 3.935
pixelwise 0.098 1.000 0.100 2.562
Pred. Sets Loc. additive 0.098 1.000 0.101 2.957
multiplicative 0.093 1.006 0.100 2.562
Cls. aps 0.072 0.994 0.061 0.993
lac 0.101 0.669 0.106 0.545
V-B CODbase
To emphasize the reproducibility and reusability of our work, we introduce a comprehensive codebase, which we refer to as the “Conformal Object Detection Codebase,”161616 https://github.com/leoandeol/cods. or CODbase for short. This open-source repository includes several widely used object detectors, including, but not limited to, YOLO and DETR.
It includes the procedures outlined in this paper, multiple loss functions, and a suite of visualization tools for clear understanding and interpretation of results. Figure 7 shows an example, where the ground truth, prediction, and corrected bounding boxes are colored green, red, and purple, respectively. The class label, top-1 predicted label, and conformal sets are also displayed.
The CODbase is implemented in Python and PyTorch, compatible with both CPU and GPU. Finally, the CODbase supports the integration of new procedures, prediction sets, and loss functions, enabling extensive benchmarks and further comparisons in future works.
V-C Benchmark & Results
The benchmark consists of a large-scale study of all combinations of the losses, prediction sets, matching functions and more (see Section V-A). This results in the largest benchmark in Conformal Object Detection, with hundreds of individual runs. Key results are presented in several tables and figures in this section, while complete results are provided in Appendix C-A.
Impact of Component Choice We begin by examining Table I. For each loss and prediction set introduced in Section IV, we present the best experimental result (i.e., the reported experiment being the best combination of the remaining parameters). It does not include a section for both classification losses, and confidence prediction sets, as only one of each is introduced in Section IV. The table is organized as follows: it is split into two, with the upper part comparing different losses, and the lower part about prediction sets. These are then subdivided into tasks, where we compare losses or prediction sets specific to each task. For instance, the first line reads: for the box_count_threshold confidence loss, and the best remaining hyperparameters (other losses, matching, etc), the confidence set size is 26.379, with an empirical test risk of 0.019. The global (joint localization and classification) empirical risk is 0.084. The experiments are conducted with DETR-50, in the $\alpha^{\mathrm{tot}}=0.1$ setting. For virtually all tasks and global risk levels, the risk is controlled, i.e., below its corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$. A few experiments showcase a risk slightly above their respective $\alpha$, likely due to randomness from sampling. We also observe that the set sizes are sufficiently small and therefore practically informative for each loss. For example, in classification, when the classes in the set are semantically close, or induce a similar response, a set size of $3$ can be considered informative. For example, a set containing pedestrian, animal and cyclist for autonomous driving. The choice of loss function significantly affects the resulting set sizes, particularly for the losses presented here. The “pixelwise” loss, a relaxation of the previous two, produces smaller corrections at the cost of only slightly weaker guarantees (e.g., a guarantee on containing most of the ground truth rather than all of it), confirming results from [15]. Similarly, the relaxed confidence loss “box_count_recall” leads to the same conclusion.
However, these losses are compared using the best possible remaining parameters. While presented set sizes are reasonably small, this is not true for all settings. Figure 6 shows the histogram of set sizes for localization and classification across numerous experiments, where each element represents a unique combination of losses, prediction sets, matching functions, and models. The histogram reveals a heavy tail: although a significant proportion of experiments produce informative set sizes (leftmost bar), a notable share results in excessively large set sizes, including bounding boxes as large as the image and conformal classification sets with dozens of classes. This implies, that for a desired error level $\alpha^{\mathrm{tot}}$, the choice of losses can lead to either informative, or overly large –unusable– sets.
Matching Functions & Error Levels We now look at Table II. It compares, for each 4 proposed matching functions in Section IV-A, and two risk levels $\alpha^{\mathrm{tot}}$, the size metrics of the three tasks, for the best performing experiment in this setting (i.e. best losses and prediction sets for each $\alpha^{\mathrm{tot}}$ and matching function). We observe that the confident set size is constant for all matching functions for $\alpha^{\mathrm{tot}}=0.1$. This is due, in our setting, to the maximum of the three risks in Eq. (4) being always the confidence risk. Indeed, when $\alpha^{\mathrm{cnf}}$ is not too high, at least a box is selected on most images and therefore by definition, the localization and classification risks in this step are null (due to $\bar{\lambda}$). However, for $\alpha^{\mathrm{tot}}=0.2$, we observe two confident set sizes, which is due to the best experiment (in the sense of localization and classification set sizes) relying on different confidence losses depending on the matching. We further observe that localization and classification set sizes are respectively minimized with the “Hausdorff” and “LAC metrics”, while “Mix” produces an interesting trade-off of the two. However, the GIoU matching, while more OD-inspired, suffers from poor performance across the board, due to the lack of connection to the subsequent losses, that use the matching. Finally, in the “Mix” setting, a large jump in classification set size can be noted. Indeed, as observed in other experiments in Appendix C-A, the impact of a risk level $\alpha$ to its associated task’s resulting set size is not smooth, it can greatly vary for a small change in target risk.
Model-Agnosticism
We now look at Table III which compares set sizes and empirical risks for DETR and YOLOv8. It is built similarly to Table I, introduced previously. We observe that for both model, the risk is controlled for all experiments, that is below the corresponding risk level $\alpha^{\mathrm{tot}}$ or $\alpha^{j}$. [^17] Using APS results in a looser bound, as discussed in [65] This result is a direct consequence of the model-agnosticism of conformal prediction and therefore of SeqCRC – it is applicable regardless of the performance or size of the model. However, the set size depends on the quality of the underlying model. Furthermore, there is a trade-off in terms of number, and size of (conformal) bounding boxes. We observe here that using YOLOv8 tends to lead to much fewer selected bounding boxes, but with much larger corrections in localization: with the same hyperparameters used for calibration, the resulting trade-off differs.
V-D Discussion
Risk Control We have observed that in all experiments, the risk is controlled (or, at worst, very close to the target), empirically verifying the guarantees presented in Theorem 2. However, it appears that the global risk guaranteed by Corollary 1 can be unnecessarily conservative, as observed in Table I and in Appendix C-A. The correction applied to obtain this global guarantee can therefore be a future point of refinement.
Guarantee strength – set size Tradeoff A shown in Table I and Table III, the further relaxed losses, as the “pixelwise” in localization, offer much smaller prediction sets. This however comes at a cost of a slightly weaker guarantee. In that case again, the two other losses provide a guarantee on the whole object being covered, while the “pixelwise”, only on the proportion of the area that is covered (i.e., missing a few pixels of a ground truth is not too punitive). We argue this approach is generally sensible and beneficial, but in some critical applications, a guarantee that the entire ground truth is contained is necessary, such as in [70].
Strengths and Limitations in Practice Turning to a more qualitative analysis, we make further observations. There are many examples of successful conformal predictions such as on Fig. 7, where the right amount of object is output, entirely contained by the slightly-corrected conformal boxes, and with a correct, singleton, classification prediction set. However, there are also shortcomings of our approach, and we focus on an example in Fig. 8. There, while the two zebras have rather satisfactory conformal predictions, there is an additional large prediction including both animals, and not linked to any ground truth. This allows us to raise two important points.
First, our approach only provides Recall guarantees, and therefore does not treat false positives as an issue. In fact, due to precision being a strongly non-monotone function of our parameters, monotonizing it while using our approach generally fails (due to the monotonized risk being too large at its lowest value).
Secondly, we have observed that the annotation in MS-COCO is very variable, and one such example, is when there are multiple occurrences of an object. Often, some individual objects are annotated, but also, there is an additional box, including all these objects, and many non annotated ones. Several such examples are presented in Appendix C-B. The model being trained on such data, repeats this pattern on some test images when not expected, and sometimes not when it is.
Finally, we recall that conformal prediction brings a guarantee on the loss of the prediction with regards to the ground truth. If the ground truth is incorrect, or unpredictable (due to inconsistencies in the annotation choices or process), and if the model does not make those errors, it will lead to arbitrarily large corrections necessary to enforce low risk on the conformal predictions, resulting in practically unusable sets.
Figure 8: Example of an artifact in conformal predictions. We observe that beyond the two correctly detected zebras, is an additional, larger, third prediction that includes both animals, but does not correspond to a ground truth. We argue this is due to artifacts in the annotation of MS-COCO, and therefore due to the pretrained model.
VI Conclusion & Future Works
In this work, we address a critical challenge in object detection: risk control. While the topic had previously been studied in some restricted cases (for a single model, or a sub-task of OD), we provide the first framework to build guaranteed prediction sets for all common tasks of object detection.
Our main contribution lies in the introduction of Sequential Conformal Risk Control (SeqCRC). Unlike previous works using CRC, SeqCRC allows to control risks of multiple dependent parameters, enabling simultaneous control of confidence thresholds, classification score and localization margin.
We then study different ways to measure prediction errors (conformal loss functions) – reviewing existing ones and proposing new ones suitable for object detection. We discuss their properties, and also consider how best to match predicted objects with the real ones (ground truths).
We test our framework through experiments, confirming that our SeqCRC method functions correctly and efficiently. A key outcome is the creation of the first benchmark for Conformal Object Detection, establishing a standardized way to compare methods in this field. Our analysis also highlights the strengths and weaknesses of this approach, identifying areas that require further research.
To support the research community, we provide an open-source codebase used to generate our benchmark. This resource includes models and procedures, enabling development of new techniques, error metrics, or guarantees based on this work.
Our analysis identifies key directions for future research. One direction is exploring less conservative approaches for obtaining global guarantees than the current one, which could lead to tighter prediction sets. Another area is the development of loss functions that better balance valuable guarantees with smaller, i.e. more informative, prediction set sizes. Extending the framework to control false positives (precision) and improving its robustness against inconsistencies in ground-truth annotations present additional challenges. Further optimization of component choices, guided by our benchmark and supported by our open-source codebase, will enhance practical applicability.
Overall, this research provides a solid foundation for building increasingly trustworthy object detection systems. We offer a novel method (SeqCRC) and valuable resources (the benchmark and codebase) to advance the development of reliable computer vision applications.
Appendix A Level of Guarantee, Aggregation & Matching
Level of Guarantee and Aggregation Functions:
On a given image, there can a priori be any number of objects (and their associated bounding boxes). On the one hand, each of those individual objects may be considered as ground truth individually. On the other hand, the ground truth can be defined as the set of all objects present on an image. Formally, the first setting corresponds to $Y_{i}=(b_{i},c_{i})$, while the latter corresponds to the setting introduced in the article, i.e. $Y_{i}=(b_{i,j},c_{i,j})_{j=1}^{n_{i}}$. Respectively, we refer to those as object-level and image-level guarantees following [14, 16, 15]. Object-level guarantees have previously been referred to as box-level in [14, 16]. However, the exact setting varied, as conformal prediction was only applied to true-positive bounding boxes (i.e., ground-truth - prediction pairs, with only pairs with IoU above a certain threshold $\beta$ considered).
Defining an object, or the set of objects on an image as the ground truth $Y_{i}$ implies a different assumption on the data generating process, and thus a different guarantee. Indeed, the expectation $\mathbb{E}\left[L_{\mathrm{test}}(\lambda)\right]$ is on the joint distribution of the calibration set and the test pair $(X_{\mathrm{test}},Y_{\mathrm{test}})$. Therefore, the guarantees hold “on average” over either image-object set pairs or over image-individual object pairs, depending on the previous choice. Importantly, the difference in definition also implies a different assumption: the pairs $(X_{i},Y_{i})$ are assumed to be independent and identically distributed. While it is beyond the scope of this work to discuss the real word validity of this assumption, the image-level one is milder than the object-level one, which implicitly assumes independence between objects of a same image.
From a practical standpoint, if an image always contains exactly one object, there is no difference. More generally, when the image-wise loss is an aggregation of object-level losses, a connection can be made between the guarantees by simple reweighing.
The choice depends on practical requirements and should be reflected in the definition of the losses. An example is a binary loss: in the object-level setting, it entails that a proportion at least $1-\alpha$ of ground truth bounding boxes are correctly covered, while in the image-level setting, it entails that, on a proportion $1-\alpha$ of images at least, all bounding boxes (or a desired proportion) are covered.
Matching:
As common in object detection [40, 43], a matching between the true bounding boxes and predicted ones is necessary, not only for evaluation here but also for calibration. We refer to by matching, a family of functions $\pi_{i}$ defined for each image $X_{i}$ that maps the index $j$ of an object on $X_{i}$ to the index of a predicted box among $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$. More formally, $\pi_{i}:\{1,\cdots,n_{i}\}\rightarrow\{1,\cdots,\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})\rvert\}$ with $n_{i}$ the number of objects on image $X_{i}$ and $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$ the sets of bounding boxes likely to correspond to objects on image $X_{i}$. We do not make further assumption on the matching algorithm. Therefore it includes injective functions, i.e. each prediction being matched to at most one ground truth. We can also consider functions where some inputs might have identic output, i.e. reusing predictions for multiple ground truth. An example of an injective matching is the commonly used Hungarian algorithm. Notably, allowing reusing a same predicted bound for more than one ground truth object is an approach to avoid the true-positive limit of [14].
Additionally, the above definition of matching fits most practical needs, but a more general approach with a one-to-many matching is also valid. Intuitively, this setting corresponds to allowing multiple predictions to be matched to a single ground truth object. That general approach includes several settings, even ones that do not explicitly consider a matching, as in [15]. Finally, prediction sets here can be reused as-is in this general setting, and losses only have to be slightly modified in most cases.
Appendix B Proofs
Throughout this section, for the sake of readability, we sometimes denote the test sample as the $(n+1)$-th sample, thus replacing $X_{\mathrm{test}},Y_{\mathrm{test}},L_{\mathrm{test}}$ by $X_{n+1},Y_{n+1},L_{n+1}$.
For readability again, we also leave out a few measure-theoretic details, such as the fact that we use the cylindrical $\sigma$-algebra (e.g., [71, p.16]) to properly define the random functions $L^{\mathrm{cnf}}_{i}$ or $L^{\bullet}_{i}$, or the fact that $\lambda^{\mathrm{cnf}}_{\pm},\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{+},\lambda^{\bullet}_{*}$ can be made measurable (i.e., that they are well-defined random variables), as well as other quantities such as $L^{\bullet}_{i}\left(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*}\right)$.
B-A Main theorem
The main theorem is reminded below. Recall that the placeholder $\bullet\in\{\mathrm{loc},\mathrm{cls}\}$ means that we work both for the localization and classification tasks, and that $\bar{\lambda}^{\mathrm{cnf}}=\max\Lambda^{\mathrm{cnf}}$.
Theorem 2.
Suppose that Assumptions 1 and 3 hold true. [^18] See Footnote 10. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
$\mathbb{E}\left[L_{\mathrm{test}}^{\bullet}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\right]\leq\alpha^{\bullet}\;.$ (7)
Under the additional assumption that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i\in\{1,\ldots,n+1\}$, we also have
$\mathbb{E}\left[L_{\mathrm{test}}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{+})\right]\leq\alpha^{\mathrm{cnf}}\;.$ (8)
Proof.
The fact that $\lambda^{\mathrm{cnf}}_{\pm}$ are well defined raises no specific issue (we use the convention $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$ in case the sets in (4) or (5) are empty). Next we start by showing that $\lambda^{\bullet}_{+}$ is well defined, and then prove the bounds (8) and (7).
Step 1: correct definition of $\lambda^{\bullet}_{+}$. We show that the set in (6) is nonempty, by distinguishing two cases. If $\lambda^{\mathrm{cnf}}_{-}=\bar{\lambda}^{\mathrm{cnf}}$, then using the assumptions $L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ and $\alpha^{\bullet}\geq\frac{B^{\bullet}}{n+1}$, we can see that the set
+=$\mathcal{L}_{+}^{\bullet}=\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{nR^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})}{n+1}+\frac{B^{\bullet}}{n+1}\leq\alpha^{\bullet}\right\}$ (18)
contains $\bar{\lambda}^{\bullet}$ and is thus nonempty. (Recall that $R^{\bullet}_{n}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})=\frac{1}{n}\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$.)
On the other hand, if $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, then by definition of $\lambda^{\mathrm{cnf}}_{-}$, [^19] In particular, when $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, the set in (5) is nonempty. by $R^{\bullet}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\bullet})\leq\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$, and by right-continuity of $R^{\bullet}_{n}$ in its first argument, we have $\frac{n}{n+1}R^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\bar{\lambda}^{\bullet})\leq\alpha^{\text{cnf}}$ almost surely. Combining this with the assumption $\alpha^{\bullet}\geq\alpha^{\text{cnf}}+\frac{B^{\bullet}}{n+1}$ implies that $\bar{\lambda}^{\bullet}\in\mathcal{L}_{+}^{\bullet}$ as in the previous case.
Therefore, in both cases, the (random) set $\mathcal{L}_{+}^{\bullet}$ is almost surely nonempty, so that $\lambda^{\bullet}_{+}=\inf\mathcal{L}_{+}^{\bullet}$ is indeed well defined.
Step 2: proof of Eq. (8). Though it almost follows from [53, Theorem 1], we briefly recall the main arguments for the sake of completeness. In this step, we make the additional assumption that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i\in\{1,\dots,n+1\}$. The proof proceeds by introducing the oracle
$\lambda^{\mathrm{cnf}}_{*}=\inf\left\{\lambda^{\mathrm{cnf}}\in\Lambda^{\mathrm{cnf}}:\tilde{R}^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}})\leq\alpha^{\text{cnf}}\right\}\;,$ (19)
where
$\tilde{R}^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}})$ $=\max\Big{\{}R^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}}),R^{\mathrm{loc}}_{n+1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}}),$
$\qquad\qquad R^{\mathrm{cls}}_{n+1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})\big{\}}\;.$ (20)
First note that $\tilde{R}^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely. (Indeed, by assumption, $R^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ and $R^{\bullet}_{n+1}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ a.s.) Therefore, the set in (19) is nonempty and $\lambda^{\mathrm{cnf}}_{*}$ is well defined. [^20] In Step 3.a below, we will also handle cases in which this set is empty. Moreover, by (19), by $R^{\mathrm{cnf}}_{n+1}\leq\tilde{R}^{\mathrm{cnf}}_{n+1}$ and by right-continuity of $R^{\mathrm{cnf}}_{n+1}$, we have $R^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}}_{*})\leq\alpha^{\text{cnf}}$ almost surely. Taking the expectation and noting that $\lambda^{\mathrm{cnf}}_{*}$ is symmetric (i.e., invariant under any permutation of the $n+1$ calibration points; see Lemma 1 in Appendix B-B below) yields $\mathbb{E}\left[L_{n+1}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{*})\right]\leq\alpha^{\mathrm{cnf}}$. The fact that $\lambda^{\mathrm{cnf}}_{+}\geq\lambda^{\mathrm{cnf}}_{*}$ (see details in Step 3.a below) and that $L_{n+1}^{\mathrm{cnf}}$ is non-increasing concludes the proof of (8).
Step 3: proof of Eq. (7). The rest of the proof is novel. Though similar in spirit to that of [53, Theorem 1], our multitask extension requires some additional care because $\lambda^{\bullet}_{+}$ uses the same calibration data as $\lambda^{\mathrm{cnf}}_{+}$. This is in contrast with, e.g., the analysis of [24, Remark 1 and Theorem 1] that works for two independent calibration datasets (data-splitting trick). Note moreover that, contrary to Step 2, in all the sequel, we do not assume that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i$.
We proceed in three steps: we first compare $\lambda^{\mathrm{cnf}}_{+}$ and $\lambda^{\bullet}_{+}$ with the oracles $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ defined in (21) and (22) below, which yields $\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\right]\leq\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]$. Then, in Step 3.c, we show that $\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]\leq\alpha^{\bullet}$.
Step 3.a: we show that $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely. To that end, recall that [^21] Though not written explicitly to simplify notation, the parameter $\lambda^{\mathrm{cnf}}$ runs over $\Lambda^{\mathrm{cnf}}$ in each of the three sets below.
$\lambda^{\mathrm{cnf}}_{+}$ $=\inf\left\{\lambda^{\mathrm{cnf}}:\frac{n\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})}{n+1}+\frac{\tilde{B}^{\text{cnf}}}{n+1}\leq\alpha^{\text{cnf}}\right\}\;,$
$\lambda^{\mathrm{cnf}}_{*}$ $=\inf\left\{\lambda^{\mathrm{cnf}}:\tilde{R}^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}})\leq\alpha^{\text{cnf}}\right\}\;,$ (21)
$\lambda^{\mathrm{cnf}}_{-}$ $=\inf\left\{\lambda^{\mathrm{cnf}}:\frac{n\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})}{n+1}+\frac{0}{n+1}\leq\alpha^{\text{cnf}}\right\}\;,$
where $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$ by convention. (The definition of $\lambda^{\mathrm{cnf}}_{*}$ in (21) is slightly more general than in (19), since empty sets are possible without the assumption on $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})$.)
Note that the three sets above, say, $A_{+}\subset A_{*}\subset A_{-}$, are nested. (This is by definition of $\tilde{R}^{\mathrm{cnf}}_{n+1}$ in (20) and because $L^{\mathrm{cnf}}_{n+1}$ and $L^{\bullet}_{n+1}$ are bounded between 00 and $\tilde{B}^{\mathrm{cnf}}=\max\{B^{\mathrm{cnf}},B^{\mathrm{loc}},B^{\mathrm{cls}}\}$ by Assumption 3.). Taking the infimum (with the convention $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$) yields $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely.
Step 3.b: we show that $\lambda^{\bullet}_{*}\leq\lambda^{\bullet}_{+}$ almost surely, where the oracle $\lambda^{\bullet}_{*}$ (which depends on the first-step oracle $\lambda^{\mathrm{cnf}}_{*}$) is defined by
$\lambda^{\bullet}_{*}=\inf\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet})\leq\alpha^{\bullet}\right\}\;.$ (22)
To see why $\lambda^{\bullet}_{*}$ is bounded from above by $\lambda^{\bullet}_{+}=\inf\mathcal{L}_{+}^{\bullet}$ (see (6) and (18)), recall that $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}$ (Step 3.a), and that the $L_{i}^{\bullet}$ are non-increasing in $\lambda^{\mathrm{cnf}}$ and bounded from above by $B^{\bullet}$ (by Assumption 3). Therefore, almost surely, for all $\lambda^{\bullet}\in\Lambda^{\bullet}$,
$\frac{1}{n+1}\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})+\frac{B^{\bullet}}{n+1}$
$\geq$ $\frac{1}{n+1}\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet})+\frac{L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet})}{n+1}\;.$
The set in (22) thus contains the set $\mathcal{L}_{+}^{\bullet}$ defined in (18), so that $\lambda^{\bullet}_{*}\leq\lambda^{\bullet}_{+}$ almost surely. Importantly, the fact that $\mathcal{L}_{+}^{\bullet}$ is a.s. nonempty also entails that the set in (22) is a.s. nonempty.
Putting Steps 3.a and 3.b together, and by monotonicity of $L_{n+1}^{\bullet}$ in both its variables, we obtain $L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\leq L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})$ a.s., so that, taking expectations,
$\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\right]\leq\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]\;.$ (23)
Step 3.c: we show that $\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]\leq\alpha^{\bullet}$, by using the fact that $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ are symmetric in the sense of Lemma 1 in Appendix B-B below.
More precisely, since the sequence $(L^{\text{cnf}}_{i},L^{\text{loc}}_{i},L^{\text{cls}}_{i})_{i=1}^{n+1}$ is i.i.d. (by Assumption 1) and thus exchangeable, and since the oracles $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ are symmetric (by Lemma 1), we have, permuting $L^{\bullet}_{n+1}$ with any $L^{\bullet}_{i}$,
$\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]$ $=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{E}\left[L^{\bullet}_{i}\left(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*}\right)\right]$
$=\mathbb{E}\left[\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}\left(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*}\right)\right]\leq\alpha^{\bullet}\;,$
where we used the fact that $\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}\left(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*}\right)\leq\alpha^{\bullet}$ almost surely. (This follows from the fact that the set in (22) is nonempty, with an infimum in $\lambda^{\bullet}_{*}$, and by right-continuity of $\lambda^{\bullet}\mapsto L^{\bullet}_{i}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet})$.) Combining (23) with the last inequality above concludes the proof of (7). ∎
B-B A simple symmetry property
[…]
Lemma 1.
The oracle parameters $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ defined in (21) and (22) are symmetric functions of $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$. That is, for any permutation $\sigma$ of $\left\{1,\ldots,n+1\right\}$, we have
$\lambda^{\mathrm{cnf}}_{*}\left((L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})\right)=$ (24)
$\lambda^{\mathrm{cnf}}_{*}\left((L^{\text{cnf}}_{\sigma(1)},L^{\text{loc}}_{\sigma(1)},L^{\text{cls}}_{\sigma(1)}),\ldots,(L^{\text{cnf}}_{\sigma(n+1)},L^{\text{loc}}_{\sigma(n+1)},L^{\text{cls}}_{\sigma(n+1)})\right)$ (25)
and
$\lambda^{\bullet}_{*}\left((L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})\right)=$ (26)
$\lambda^{\bullet}_{*}\left((L^{\text{cnf}}_{\sigma(1)},L^{\text{loc}}_{\sigma(1)},L^{\text{cls}}_{\sigma(1)}),\ldots,(L^{\text{cnf}}_{\sigma(n+1)},L^{\text{loc}}_{\sigma(n+1)},L^{\text{cls}}_{\sigma(n+1)})\right).$ (27)
Proof.
The proof is rather straightforward, but we provide it for the sake of completeness. We start by showing that $\lambda^{\mathrm{cnf}}_{*}$ is symmetric. Recall from (21) that
$\lambda^{\mathrm{cnf}}_{*}=\inf\left\{\lambda^{\mathrm{cnf}}\in\Lambda^{\mathrm{cnf}}:\tilde{R}^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}})\leq\alpha^{\text{cnf}}\right\}$
(with the convention $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$), and note from (20) that $\tilde{R}^{\mathrm{cnf}}_{n+1}$ is a maximum of three terms that are invariant under any permutation of the $3$-tuples $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$. Therefore, $\lambda^{\mathrm{cnf}}_{*}$ is symmetric.
We now prove the symmetry of $\lambda^{\bullet}_{*}$, which is only a little more subtle. Indeed, note from (22) that the definition
$\lambda^{\bullet}_{*}=\inf\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet})\leq\alpha^{\bullet}\right\}$
involves the oracle parameter $\lambda^{\mathrm{cnf}}_{*}$. Since both the sum operator and $\lambda^{\mathrm{cnf}}_{*}$ (by the first part of the proof) are permutation-invariant, so is $\lambda^{\bullet}_{*}$. This concludes the proof. ∎
(a) Successful detection of bears
(b) Successful detection of a clock tower
(c) Technically successful detection of an airplane
Figure 9: Examples of two successful conformal predictions, and one technically successful: this is a consequence of our recall-based guarantee.
Appendix C Details on Experiments
In this section, we introduce the complete benchmark of our experiments, as well as additional experiments which include discussions on metrics in OD and conformal OD, localization margins and monotonicity.
C-A Complete benchmark
We include the complete results of our experiments here in Tables V, VI, VII, and VIII. We now introduce multiple additional visualizations of our best-performing experiment run, for additional discussions on the limitations of conformal object detection. On Fig. 9, we observe two successful detections, with singleton classification prediction sets and small localization margins. However, on the third image, the moon is predicted to be a kite. While this is incorrect, due to the monotonicity constraints of CRC, we may only provide recall guarantees, i.e., ensure that we do not miss the ground truth. In that framework, additional predictions such as the moon as a kite here, are not penalized. In Fig. 10, we observe another instance of the same mistake, however, here the boat is detected twice, once excluding, and once including its mast. We however, observe two failure cases of our guarantee, corresponding to previously discussed issues. Indeed, in both examples, there are abuses of labeling: on one books are sometimes annotated individually, and sometimes as a block. On the second, a statue is annotated as a person. The variability in the definition of ground truths leads to perceived failures, that are not necessarily actual failures. As visible on both images, the clock and statue (person) are respectively assigned a book and a clock predictions by our matching. In order to reach lower levels of error rates $\alpha$, these mistakes would have to be controlled, and would result in a need for significant margins at inference (or models that are able to learn this bias).
C-B Additional Experiments
In this section, we discuss three additional experiments: one on the non-monotonicity of losses, one on the correlation between OD and CP metrics, the other on the number of margins used for conformal predicted boxes.
C-B1 Monotonicty
In this section, we conduct an additional study on the monotonicity of loss functions, used or considered at a point in our work.
C-B2 mAP vs CP
In the course of our exploratory experiments, we encountered an intriguing yet logically consistent phenomenon: the performance metrics commonly used in Object Detection (OD) do not necessarily align with the efficacy of Conformal Prediction (CP) techniques, particularly in terms of set size. This is exemplified in Figure 11.
To clarify, consider two hypothetical models, Model 1 and Model 2, with Model 2 demonstrating superior performance in terms of mean Average Precision (mAP). This superiority is evidenced by its larger Intersection over Union (IoU) values with the ground truth across four sampled images, attributed in part to its tendency to predict larger bounding boxes. Notably, these models exhibit distinct error characteristics: Model 1 primarily errs in scaling, whereas Model 2’s errors are predominantly cause by translations. However, a critical insight emerges when considering the necessity to ensure that the ground truth is entirely captured within the predicted prediction’s boundaries. Both models require the addition of a comparable margin to their predictions to meet this criterion. Consequently, due to Model 2’s initially larger predictions, the resultant conformal boxes are significantly larger than those of Model 1.
In a sense, a model that fails in a consistent and predictable manner (“reliably”) can be more efficiently adjusted and at a lower “cost” in CP terms, as opposed to a model that, despite its higher performance metrics, fails in an unpredictable or “random’ fashion.
(a) Technically successful detection of a boat
(b) Unsuccessful prediction of a clock
(c) Unsuccessful prediction of a statue (person)
Figure 10: Examples of one technically successful conformal prediction (as in Fig. 9) and two failures, due to a missing prediction.
C-B3 Margins
In our research, we executed a targeted experiment to ascertain the optimal strategy for applying margins to the predicted bounding boxes in Object Detection. Our primary objective was to determine whether uniform margins should be applied across all four sides of a bounding box, whether separate margins should be used for height and width, or whether distinct margins should be applied to each side. It is important to note that while intuitively, the latter approach may seem advantageous in most scenarios, it inherently equates to conducting a multiple-hypothesis test, necessitating a correction for multiple comparisons.
TABLE IV: Comparison of Margins in pixels, Coverages and Set sizes for considering {1,2,4} distinct values for margin corrections (with Bonferroni correction), for $\alpha=0.1$ (resp. $\alpha=0.05$, $\alpha=0.025$ with correction).
N. margins Value of the margins Cov Set Size
Left Top Right Bottom
1 11.88 11.88 11.88 11.88 96.30 % 142
2 19.58 16.18 19.58 16.18 97.43 % 145
4 26.34 24.89 28.11 14.30 97.99 % 151
In our analysis, we employed a Bonferroni correction to account for this multiplicity. However, the application of such a correction incurs a certain “cost” and may not always yield superior results. For instance, as evidenced in Table IV of our study, we observed a trend wherein the more margins incorporated, the more conservative our procedure became. Consequently, in our specific case, it was more efficient to apply a single uniform margin in all directions.
It is crucial to underscore that this conclusion is not universally applicable across all datasets and models. For example, in a dataset where objects predominantly align along the same row, a more prudent approach might involve applying larger corrections horizontally and smaller vertically. This experimental approach is not resource-intensive and can be readily replicated on different datasets using our publicly available repository.
Appendix D Effect of Object Detection and Conformal Settings
Obtaining statistically valid predictions in Object Detection (OD) is a multifaceted problem, dependent upon the desired level of guarantee and the specific configuration of object detectors. A pivotal issue concerns the ambiguity of the concept of “ground truth”. This can be conceptualized as either a set of bounding boxes for an image or as individual bounding boxes.
In this paper, we adopt the latter approach, the box-level setting, for its clarity in comparison to the set-based approach.
Our study also touches upon the levels of guarantees in terms of confidence, localization, and classification within OD. Notably, classification and localization are inherently box-level concepts. Classification presupposes the correct identification of an object within a predicted bounding box, rendering predictions meaningless in the absence of an actual object. In our Conformal Prediction (CP) approach, we primarily control for the classification accuracy of the predicted bounding box nearest to the ground truth, not taking into account the classification accuracy of other predicted boxes that might have more accurately identified the object, albeit with less precise coordinates. Regarding confidence thresholds, formalization is challenging due to it not actually being a ground truth. In this context, we opted for an image-level guarantee, ensuring that, with a probability of at least $1-\alpha$, the number of predictions matches or exceeds the number of ground truths, irrespective of their spatial accuracy.
Lastly, regarding the $x_{1}y_{1}x_{2}y_{2}$ (coordinates of two corners) versus $xywh$ (coordinates of the center + width and height) bounding box representations, while fundamentally equivalent, their application within CP algorithms yields divergent results. The $xywh$ format inherently involves symmetric corrections and a translational adjustment, whereas the $xyxy$ format necessitates independent corrections for each side. However, these approaches can be reformulated to yield equivalent predictive sets.
Appendix E Additional Losses
We here discuss losses not mentioned in the article, and generally not tested. They are presented for information, and potential future works when appropriate needs arise.
As discussed in details in Appendix A, many natural image-level losses are aggregations of object-level losses. While there are exceptions to that, the losses presented in this appendix are of this form. Formally, denote by $\Phi:S^{\infty}(\mathbb{R})\rightarrow\mathbb{R}$ an aggregation function, mapping a sequence $S$ of arbitrary length – with $S_{i}$ valued in $\mathbb{R}$ – to $\mathbb{R}$. We introduce three main aggregation strategies.
  1. 1.
    Average: $\Phi_{\mathrm{avg}}(S)=\frac{1}{\lvert S\rvert}\sum_{i=1}^{\lvert S\rvert}S_{i}$
  2. 2.
    Max: $\Phi_{\mathrm{max}}(S)=\max_{i}S_{i}$
  3. 3.
    Thresholded Average (at level $\tau$): $\Phi_{\mathrm{thr}}(S)=\mathbbm{1}_{\Phi_{\mathrm{thr}}(S)>\tau}$
In the next sections, we therefore introduce object-level losses, than can be used with any of the aforementioned aggregation functions.
E-A Confidence
The losses introduced in our work for confidence focus on the number of prediction, regardless of their location on the image. It is common for most confident boxes to be located close to true objects with well-performing predictors. We have indeed observed that accounting for the localization of the prediction at the confidence step was not necessary with our models and dataset. This may however, not be the case generally, and we therefore introduce the following losses, which account only for prediction sufficiently close to the ground truth. These losses are presented for an image $x$, for a true box $b$, and the set of conformal bounding boxes $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$. First, one can simply consider the inclusion of the ground truth:
$L^{\mathrm{cnf}}_{\mathrm{incl}}(\lambda^{\mathrm{cnf}})=\begin{cases}0&\text{if }\exists\hat{b}\in\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x),b\subseteq\hat{b}\\ 1&\text{ otherwise}.\end{cases}$
This criterion is however very stringent and can result very quickly in a very low confidence threshold, i.e., numerous predictions being kept. We therefore recommend to use it only after applying a scaling or margin on the predictions $\hat{b}$ by a factor decided a priori (hyperparameter).
Figure 11: Higher CP performance (small correction prescribed) does not necessarily correlate to a higher mean Average Precision (mAP). Considering a simple additive margin in CP, an always-smaller but always-centered prediction performs worse in mAP, but requires significantly less conformal correction than a better mAP-performing model, which often overshoots the ground truth box, uniformly at random in each direction.
An alternative is to replace the inclusion criterion $\subseteq$ by a distance, formalized by
$L^{\mathrm{cnf}}_{\mathrm{dst}}(\lambda^{\mathrm{cnf}})=\begin{cases}0&\text{ if }\exists\hat{b}\in\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x),d((b,c),(\hat{b},\hat{c}))<\tau\\ 1&\text{ otherwise},\end{cases}$
where $d:(\mathcal{B}\times\mathcal{C})\times(\mathcal{B}\times\Sigma_{K-1})\rightarrow\mathbb{R}^{+}$ is any function measuring a distance (particularly, it does not need to be symmetric), and $\tau$ is a pre-defined threshold for a prediction to be considered sufficiently close to a ground truth. Note that we define the loss to be 00 when there are no ground truths on an image. Note that the labels $c$ and $\hat{c}$ are also included, and may or not be used (such as to set arbitrarily high values to significantly different classes). We recommend for the sake of consistency to use the same distance as picked for matching, discussed in Section IV-A.
An alternative approach is to the average distance of the predictions from the ground truths.
$L^{\mathrm{cnf}}_{\mathrm{dist}}(\lambda^{\mathrm{cnf}})=\min_{(\hat{b},\hat{c})\in\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)}d\left((b,c),(\hat{b},\hat{c})\right),$
where $d:(\mathcal{B}\times\mathcal{C})\times(\mathcal{B}\times\Sigma_{K-1})\rightarrow[0,1]$ is any normalized distance (or divergence). Again, we recommend to use the same distance used for matching predictions and ground truths, in order to obtain more interpretable guarantees.
Nevertheless, an issue pertains: the prediction sets for confidence depend strongly on the confidence score and are not adaptive. Further work in this direction, such as building more complex, uncertainty-aware confidence scores, or combining them is warranted.
E-B Localization
The prediction sets introduced in the main article are simple, but suffer from a lack of adaptivity. This is partly due to the lack of direct notion of uncertainty for predicted bounding boxes. A recent work, BayesOD [8] method replaces Non-Maximum Suppression by a Bayesian fusion, which results in uncertainty metrics on the coordinates of the corners of the predictions. Such information can be replace the width/height estimates in the multiplicative correction. We expect the resulting conformal boxes could grant better adaptivity:
$\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}(x)_{k}=\hat{b}_{k}+\begin{pmatrix}\hat{\sigma}^{1}_{k}&\hat{\sigma}^{2}_{k}&\hat{\sigma}^{3}_{k}&\hat{\sigma}^{4}_{k}\end{pmatrix}\,,$ (28)
where the $\hat{\sigma}^{i}_{k}$ would quantify uncertainty of the $k$-th box coordinate, such as empirical standard deviations.
E-C Classification
Classification is, unlike the two preceding ones, a fairly common task in conformal prediction. We therefore refer the reader to the large literature on the topic, including but not limited to [65, 66, 67, 68].
E-D Note on Monotonicity
While we mention several times throughout the article the necessity of the loss to be monotone in all its parameters, it is in fact not strictly necessary. Since our optimization algorithm, introduced in the next section, allows to control risks that are non-monotone by monotonizing them on the fly, any losses can be used. However, at the two extremes lie opposite results. On the one hand, when monotonicity is verified, our algorithm offers no penalty. On the other hand, when the loss is strongly non-monotone (think $x\rightarrow x^{2}$), the resulting prediction sets will be overly large and non-informative.
Additionally, we could in principle enforce monotonicity in $\lambda^{\mathrm{cnf}}$ by computing a non-injective matching $\pi_{x}$ with a distance $d$ that mimics a given localization loss (so that adding boxes in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ could only decrease this loss). However, this would not fix the lack of monotonicity for classification losses (see Remark 2), since we use the same matching $\pi_{x}$ for the two tasks.
Appendix F Algorithms
In this section, we introduce two algorithms used for optimization of the first- and second-step parameters in our proposed SeqCRC.
Algorithm 3 SeqCRC Monotonizing Optimization for Step 1 (Confidence)
1:Loss functions $L^{\mathrm{cnf}}_{i}$, $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$, error level $\alpha^{\mathrm{cnf}}$, Confidence scores $C$
2:Confidence Parameter $\lambda^{\mathrm{cnf}}$
3:Flatten the set of sets of confidence scores $C$ into a single array
4:Create an array of same size, with values the index of image of origin of the associated confidence score in the previous array
5:Sort simultaneously both arrays by decreasing order of confidence scores, denoted by $\mathcal{C}$ and $\mathcal{I}$ respectively.
6:$\lambda^{\mathrm{cnf}}\leftarrow C_{0}$
7:Define a loss array $L$
8:for $i\in 1,\dots,n$ do
9:     Update losses with $\lambda^{\mathrm{cnf}}$ : $L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$, $L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$
10:end for
11:Compute risks with updated losses $R^{\mathrm{cnf}}=\sum_{i=1}^{n}L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $R^{\mathrm{loc}}=\sum_{i=1}^{n}L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$ and $R^{\mathrm{cls}}=\sum_{i=1}^{n}L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$,
12:Compute the maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$
13:for $j\in 2,\dots,\lvert\mathcal{C}\rvert$ do
14:     Update previous risk $R^{\prime}=R$
15:     Update previous $\lambda^{\mathrm{cnf}}{}^{\prime}=\lambda^{\mathrm{cnf}}$
16:     Get current confidence score $c\leftarrow\mathcal{C}(j)$
17:     Get index of the score of the previous image $i\leftarrow\mathcal{I}(j-1)$
18:     Set $\lambda^{\mathrm{cnf}}\leftarrow 1-c$
19:     Update the losses of previous image with new $\lambda^{\mathrm{cnf}}$ : $L^{\mathrm{cnf}}_{i-1}(\lambda^{\mathrm{cnf}})$, $L^{\mathrm{loc}}_{i-1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$, $L^{\mathrm{cls}}_{i-1}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{cls}})$
20:     Recompute risks with updated losses $R^{\mathrm{cnf}}=\sum_{i=1}^{n}L^{\mathrm{cnf}}_{i}(\lambda^{\mathrm{cnf}})$, $R^{\mathrm{loc}}=\sum_{i=1}^{n}L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$ and $R^{\mathrm{cls}}=\sum_{i=1}^{n}L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\mathrm{loc}})$,
21:     Compute the maximum risk $R=\max\{R^{\mathrm{cnf}},R^{\mathrm{loc}},R^{\mathrm{cls}}\}$
22:     if $R<R^{\prime}$ then
23:         $R\leftarrow R^{\prime}$
24:     end if
25:Stopping condition:
26:     if $\frac{n}{n+1}R+\frac{B}{n+1}>\alpha^{\mathrm{cnf}}$ then return previous $\lambda^{\mathrm{cnf}}{}^{\prime}$
27:     end if
28:end for
Algorithm 4 SeqCRC Monotonizing Optimization for Step 2 (Localization/Classification)
1:Loss functions $L^{\bullet}_{i}$, error level $\alpha^{\bullet}$, Confidence scores $C$, a lower $l$ and upper $u$ bound for $\lambda^{\bullet}$, a number of steps $S$, a lower confidence parameter $\lambda^{\mathrm{cnf}}_{-}$
2:Create an empty array $\hat{\Lambda}$
3:for $k\in 1,\cdot,S$ do
4:     $\lambda^{\bullet}\leftarrow\frac{l+u}{2}$
5:     Flatten the set of sets of confidence scores $C$ into a single array
6:     Create an array of same size, with values the index of image of origin of the associated confidence score in the previous array
7:     Sort simultaneously both arrays by decreasing order of confidence scores, denoted by $\mathcal{C}$ and $\mathcal{I}$ respectively.
8:     $\lambda^{\mathrm{cnf}}\leftarrow C_{0}$
9:     Define a loss array $L$
10:     for $i\in 1,\dots,n$ do
11:         Update $\Lambda^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
12:     end for
13:     Compute risk $R=\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}})$
14:     for $j\in 2,\dots,\lvert\mathcal{C}\rvert$ do
15:         Update previous risk $R^{\prime}=R$
16:         Update previous $\lambda^{\mathrm{cnf}}{}^{\prime}=\lambda^{\mathrm{cnf}}$
17:         Get current confidence score $c\leftarrow\mathcal{C}(j)$
18:         Get index of the score of the previous image $i\leftarrow\mathcal{I}(j-1)$
19:         Set $\lambda^{\mathrm{cnf}}\leftarrow 1-c$
20:         Update the loss of previous image with new $\lambda^{\mathrm{cnf}}$ : $L^{\bullet}_{i-1}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
21:         Recompute risk with updated loss $R=\sum_{i=1}^{n}L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$
22:         if $R<R^{\prime}$ then
23:              $R\leftarrow R^{\prime}$
24:         end ifStopping condition:
25:         if $\lambda^{\mathrm{cnf}}=\lambda^{\mathrm{cnf}}_{-}$ then
26:              break
27:         end if
28:     end for
29:     if $\frac{n}{n+1}R+\frac{B}{n+1}\leq\alpha^{\bullet}$ then
30:         Append $\lambda^{\bullet}$ to $\hat{\Lambda}$
31:         $u\leftarrow\lambda^{\bullet}$
32:     else
33:         $l\leftarrow\lambda^{\bullet}$
34:     end ifreturn The last element of $\hat{\Lambda}$
35:end for
TABLE V: Experimental Results using the DETR-50 model for $\alpha_{tot}=0.1$
Matching Function Conf. Loss Loc. Loss Cls. Pred. Set Loc. Pred. Set. Confidence Localization Classification Global
Size Risk Size Risk Size Risk Risk
mix box_count_threshold thresholded lac additive 26.3792 0.0188 4.5749 0.0460 17.3165 0.0458 0.0857
mix box_count_threshold thresholded lac multiplicative 26.3792 0.0188 2.5006 0.0516 17.3165 0.0458 0.0946
mix box_count_threshold thresholded aps additive 26.3792 0.0188 4.5749 0.0440 23.4177 0.0415 0.0788
mix box_count_threshold thresholded aps multiplicative 26.3792 0.0188 2.1484 0.0484 16.2460 0.0491 0.0939
mix box_count_threshold pixelwise lac additive 26.3792 0.0188 1.5091 0.0484 13.4423 0.0476 0.0850
mix box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 1.0047 0.0510 13.4423 0.0476 0.0942
mix box_count_threshold pixelwise aps additive 26.3792 0.0188 1.5091 0.0484 14.3149 0.0484 0.0866
mix box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 1.0024 0.0510 12.3572 0.0484 0.0933
mix box_count_threshold boxwise lac additive 26.3792 0.0188 1.8151 0.0484 12.4177 0.0484 0.0860
mix box_count_threshold boxwise lac multiplicative 26.3792 0.0188 1.2427 0.0510 12.4177 0.0484 0.0942
mix box_count_threshold boxwise aps additive 26.3792 0.0188 1.8151 0.0484 13.7468 0.0484 0.0876
mix box_count_threshold boxwise aps multiplicative 26.3792 0.0188 1.2427 0.0510 12.4843 0.0484 0.0942
mix box_count_recall thresholded lac additive 18.8680 0.0186 4.5749 0.0460 17.3165 0.0458 0.0857
mix box_count_recall thresholded lac multiplicative 18.8680 0.0186 2.5006 0.0516 17.3165 0.0458 0.0946
mix box_count_recall thresholded aps additive 18.8680 0.0186 4.5749 0.0440 23.4177 0.0415 0.0788
mix box_count_recall thresholded aps multiplicative 18.8680 0.0186 2.1484 0.0484 16.2460 0.0491 0.0939
mix box_count_recall pixelwise lac additive 18.8680 0.0186 1.5091 0.0484 13.4423 0.0476 0.0850
mix box_count_recall pixelwise lac multiplicative 18.8680 0.0186 1.0047 0.0510 13.4423 0.0476 0.0942
mix box_count_recall pixelwise aps additive 18.8680 0.0186 1.5091 0.0484 14.3149 0.0484 0.0866
mix box_count_recall pixelwise aps multiplicative 18.8680 0.0186 1.0024 0.0510 12.3572 0.0484 0.0933
mix box_count_recall boxwise lac additive 18.8680 0.0186 1.8151 0.0484 12.4177 0.0484 0.0860
mix box_count_recall boxwise lac multiplicative 18.8680 0.0186 1.2427 0.0510 12.4177 0.0484 0.0942
mix box_count_recall boxwise aps additive 18.8680 0.0186 1.8151 0.0484 13.7468 0.0484 0.0876
mix box_count_recall boxwise aps multiplicative 18.8680 0.0186 1.2427 0.0510 12.4843 0.0484 0.0942
giou box_count_threshold thresholded lac additive 26.3792 0.0188 35.0661 0.0451 47.6759 0.0486 0.0879
giou box_count_threshold thresholded lac multiplicative 26.3792 0.0188 29.1481 0.0502 47.6759 0.0486 0.0940
giou box_count_threshold thresholded aps additive 26.3792 0.0188 35.0661 0.0443 46.7573 0.0460 0.0840
giou box_count_threshold thresholded aps multiplicative 26.3792 0.0188 28.3221 0.0490 45.7192 0.0460 0.0903
giou box_count_threshold pixelwise lac additive 26.3792 0.0188 30.3110 0.0476 47.9874 0.0484 0.0888
giou box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 27.6434 0.0499 47.9874 0.0484 0.0930
giou box_count_threshold pixelwise aps additive 26.3792 0.0188 30.3110 0.0476 46.7573 0.0460 0.0862
giou box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 27.4804 0.0490 45.7192 0.0460 0.0903
giou box_count_threshold boxwise lac additive 26.3792 0.0188 36.2020 0.0476 46.8872 0.0476 0.0871
giou box_count_threshold boxwise lac multiplicative 26.3792 0.0188 30.1844 0.0499 46.8872 0.0476 0.0923
giou box_count_threshold boxwise aps additive 26.3792 0.0188 36.2020 0.0476 47.0217 0.0460 0.0855
giou box_count_threshold boxwise aps multiplicative 26.3792 0.0188 30.0534 0.0490 45.7192 0.0460 0.0903
giou box_count_recall thresholded lac additive 18.8680 0.0186 35.0661 0.0451 47.6759 0.0486 0.0879
giou box_count_recall thresholded lac multiplicative 18.8680 0.0186 29.1481 0.0502 47.6759 0.0486 0.0940
giou box_count_recall thresholded aps additive 18.8680 0.0186 35.0661 0.0443 46.7573 0.0460 0.0840
giou box_count_recall thresholded aps multiplicative 18.8680 0.0186 28.3221 0.0490 45.7192 0.0460 0.0903
giou box_count_recall pixelwise lac additive 18.8680 0.0186 30.3110 0.0476 47.9874 0.0484 0.0888
giou box_count_recall pixelwise lac multiplicative 18.8680 0.0186 27.6434 0.0499 47.9874 0.0484 0.0930
giou box_count_recall pixelwise aps additive 18.8680 0.0186 30.3110 0.0476 46.7573 0.0460 0.0862
giou box_count_recall pixelwise aps multiplicative 18.8680 0.0186 27.4804 0.0490 45.7192 0.0460 0.0903
giou box_count_recall boxwise lac additive 18.8680 0.0186 36.2020 0.0476 46.8872 0.0476 0.0871
giou box_count_recall boxwise lac multiplicative 18.8680 0.0186 30.1844 0.0499 46.8872 0.0476 0.0923
giou box_count_recall boxwise aps additive 18.8680 0.0186 36.2020 0.0476 47.0217 0.0460 0.0855
giou box_count_recall boxwise aps multiplicative 18.8680 0.0186 30.0534 0.0490 45.7192 0.0460 0.0903
hausdorff box_count_threshold thresholded lac additive 26.3792 0.0188 1.6257 0.0452 44.6030 0.0503 0.0966
hausdorff box_count_threshold thresholded lac multiplicative 26.3792 0.0188 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_threshold thresholded aps additive 26.3792 0.0188 1.6257 0.0444 44.2706 0.0476 0.0849
hausdorff box_count_threshold thresholded aps multiplicative 26.3792 0.0188 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_threshold pixelwise lac additive 26.3792 0.0188 1.0936 0.0476 44.6030 0.0503 0.0987
hausdorff box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 1.0006 0.0508 44.6030 0.0503 0.1000
hausdorff box_count_threshold pixelwise aps additive 26.3792 0.0188 1.0936 0.0476 44.2706 0.0476 0.0853
hausdorff box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 1.0006 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_threshold boxwise lac additive 26.3792 0.0188 1.5783 0.0491 44.6030 0.0503 0.0987
hausdorff box_count_threshold boxwise lac multiplicative 26.3792 0.0188 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_threshold boxwise aps additive 26.3792 0.0188 1.5783 0.0491 44.2706 0.0476 0.0853
hausdorff box_count_threshold boxwise aps multiplicative 26.3792 0.0188 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall thresholded lac additive 18.8680 0.0186 1.6257 0.0452 44.6030 0.0503 0.0966
hausdorff box_count_recall thresholded lac multiplicative 18.8680 0.0186 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_recall thresholded aps additive 18.8680 0.0186 1.6257 0.0444 44.2706 0.0476 0.0849
hausdorff box_count_recall thresholded aps multiplicative 18.8680 0.0186 1.0562 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall pixelwise lac additive 18.8680 0.0186 1.0936 0.0476 44.6030 0.0503 0.0987
hausdorff box_count_recall pixelwise lac multiplicative 18.8680 0.0186 1.0006 0.0508 44.6030 0.0503 0.1000
hausdorff box_count_recall pixelwise aps additive 18.8680 0.0186 1.0936 0.0476 44.2706 0.0476 0.0853
hausdorff box_count_recall pixelwise aps multiplicative 18.8680 0.0186 1.0006 0.0499 44.2706 0.0476 0.0915
hausdorff box_count_recall boxwise lac additive 18.8680 0.0186 1.5783 0.0491 44.6030 0.0503 0.0987
hausdorff box_count_recall boxwise lac multiplicative 18.8680 0.0186 1.0562 0.0509 44.6030 0.0503 0.1000
hausdorff box_count_recall boxwise aps additive 18.8680 0.0186 1.5783 0.0491 44.2706 0.0476 0.0853
hausdorff box_count_recall boxwise aps multiplicative 18.8680 0.0186 1.0562 0.0499 44.2706 0.0476 0.0915
lac box_count_threshold thresholded lac additive 26.3792 0.0188 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_threshold thresholded lac multiplicative 26.3792 0.0188 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_threshold thresholded aps additive 26.3792 0.0188 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_threshold thresholded aps multiplicative 26.3792 0.0188 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_threshold pixelwise lac additive 26.3792 0.0188 17.3935 0.0491 1.0651 0.0503 0.0834
lac box_count_threshold pixelwise lac multiplicative 26.3792 0.0188 13.0506 0.0509 1.0651 0.0503 0.0909
lac box_count_threshold pixelwise aps additive 26.3792 0.0188 17.3935 0.0491 1.0440 0.0484 0.0819
lac box_count_threshold pixelwise aps multiplicative 26.3792 0.0188 13.0506 0.0509 1.0440 0.0484 0.0909
lac box_count_threshold boxwise lac additive 26.3792 0.0188 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_threshold boxwise lac multiplicative 26.3792 0.0188 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_threshold boxwise aps additive 26.3792 0.0188 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_threshold boxwise aps multiplicative 26.3792 0.0188 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_recall thresholded lac additive 18.8680 0.0186 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_recall thresholded lac multiplicative 18.8680 0.0186 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_recall thresholded aps additive 18.8680 0.0186 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_recall thresholded aps multiplicative 18.8680 0.0186 15.2387 0.0499 1.0440 0.0484 0.0909
lac box_count_recall pixelwise lac additive 18.8680 0.0186 17.3935 0.0491 1.0651 0.0503 0.0834
lac box_count_recall pixelwise lac multiplicative 18.8680 0.0186 13.0506 0.0509 1.0651 0.0503 0.0909
lac box_count_recall pixelwise aps additive 18.8680 0.0186 17.3935 0.0491 1.0440 0.0484 0.0819
lac box_count_recall pixelwise aps multiplicative 18.8680 0.0186 13.0506 0.0509 1.0440 0.0484 0.0909
lac box_count_recall boxwise lac additive 18.8680 0.0186 19.5160 0.0476 1.0651 0.0503 0.0828
lac box_count_recall boxwise lac multiplicative 18.8680 0.0186 15.2387 0.0499 1.0651 0.0503 0.0909
lac box_count_recall boxwise aps additive 18.8680 0.0186 19.5160 0.0476 1.0440 0.0484 0.0819
lac box_count_recall boxwise aps multiplicative 18.8680 0.0186 15.2387 0.0499 1.0440 0.0484 0.0909
TABLE VI: Experimental Results using the DETR-50 model for $\alpha_{tot}=0.2$
Matching Function Conf. Loss Loc. Loss Cls. Pred. Set Loc. Pred. Set. Confidence Localization Classification Global
Size Risk Size Risk Size Risk Risk
mix box_count_threshold thresholded lac additive 23.9744 0.0272 2.9348 0.0940 12.9752 0.0968 0.1728
mix box_count_threshold thresholded lac multiplicative 23.9744 0.0272 1.8538 0.1016 12.9752 0.0968 0.1830
mix box_count_threshold thresholded aps additive 23.9744 0.0272 2.9348 0.0924 10.6293 0.0945 0.1618
mix box_count_threshold thresholded aps multiplicative 23.9744 0.0272 1.6141 0.0984 10.2883 0.0953 0.1797
mix box_count_threshold pixelwise lac additive 23.9744 0.0272 1.0432 0.0984 0.9060 0.1000 0.1736
mix box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 0.9998 0.1016 0.9060 0.1000 0.1830
mix box_count_threshold pixelwise aps additive 23.9744 0.0272 1.0432 0.0984 0.9937 0.0724 0.1500
mix box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 0.9998 0.1016 0.9772 0.0724 0.1541
mix box_count_threshold boxwise lac additive 23.9744 0.0272 1.3962 0.0984 0.9060 0.1000 0.1736
mix box_count_threshold boxwise lac multiplicative 23.9744 0.0272 1.0432 0.1016 0.9060 0.1000 0.1830
mix box_count_threshold boxwise aps additive 23.9744 0.0272 1.3962 0.0984 0.9937 0.0724 0.1500
mix box_count_threshold boxwise aps multiplicative 23.9744 0.0272 1.0432 0.1016 0.9772 0.0724 0.1541
mix box_count_recall thresholded lac additive 14.4056 0.0262 2.9348 0.0940 12.9752 0.0968 0.1728
mix box_count_recall thresholded lac multiplicative 14.4056 0.0262 1.8538 0.1016 12.9752 0.0968 0.1830
mix box_count_recall thresholded aps additive 14.4056 0.0262 2.9348 0.0924 10.6293 0.0945 0.1618
mix box_count_recall thresholded aps multiplicative 14.4056 0.0262 1.6141 0.0984 10.2883 0.0953 0.1797
mix box_count_recall pixelwise lac additive 14.4056 0.0262 1.0432 0.0984 0.9060 0.1000 0.1736
mix box_count_recall pixelwise lac multiplicative 14.4056 0.0262 0.9998 0.1016 0.9060 0.1000 0.1830
mix box_count_recall pixelwise aps additive 14.4056 0.0262 1.0432 0.0984 0.9937 0.0724 0.1500
mix box_count_recall pixelwise aps multiplicative 14.4056 0.0262 0.9998 0.1016 0.9772 0.0724 0.1541
mix box_count_recall boxwise lac additive 14.4056 0.0262 1.3962 0.0984 0.9060 0.1000 0.1736
mix box_count_recall boxwise lac multiplicative 14.4056 0.0262 1.0432 0.1016 0.9060 0.1000 0.1830
mix box_count_recall boxwise aps additive 14.4056 0.0262 1.3962 0.0984 0.9937 0.0724 0.1500
mix box_count_recall boxwise aps multiplicative 14.4056 0.0262 1.0432 0.1016 0.9772 0.0724 0.1541
giou box_count_threshold thresholded lac additive 23.9744 0.0272 28.1254 0.0932 36.1885 0.0976 0.1708
giou box_count_threshold thresholded lac multiplicative 23.9744 0.0272 24.4546 0.1000 36.1885 0.0976 0.1830
giou box_count_threshold thresholded aps additive 23.9744 0.0272 28.1254 0.0924 34.8062 0.0953 0.1602
giou box_count_threshold thresholded aps multiplicative 23.9744 0.0272 23.8266 0.0984 35.1776 0.0968 0.1805
giou box_count_threshold pixelwise lac additive 23.9744 0.0272 25.1563 0.0952 35.0695 0.1000 0.1716
giou box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 24.2717 0.1000 35.0695 0.1000 0.1830
giou box_count_threshold pixelwise aps additive 23.9744 0.0272 25.1563 0.0952 34.8062 0.0953 0.1602
giou box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 23.9725 0.0984 35.1776 0.0968 0.1805
giou box_count_threshold boxwise lac additive 23.9744 0.0272 28.6734 0.0952 35.4345 0.0992 0.1716
giou box_count_threshold boxwise lac multiplicative 23.9744 0.0272 25.0144 0.1000 35.4345 0.0992 0.1830
giou box_count_threshold boxwise aps additive 23.9744 0.0272 28.6734 0.0952 34.8062 0.0953 0.1602
giou box_count_threshold boxwise aps multiplicative 23.9744 0.0272 24.8396 0.0984 35.1776 0.0968 0.1805
giou box_count_recall thresholded lac additive 14.4056 0.0262 28.1254 0.0932 36.1885 0.0976 0.1708
giou box_count_recall thresholded lac multiplicative 14.4056 0.0262 24.4546 0.1000 36.1885 0.0976 0.1830
giou box_count_recall thresholded aps additive 14.4056 0.0262 28.1254 0.0924 34.8062 0.0953 0.1602
giou box_count_recall thresholded aps multiplicative 14.4056 0.0262 23.8266 0.0984 35.1776 0.0968 0.1805
giou box_count_recall pixelwise lac additive 14.4056 0.0262 25.1563 0.0952 35.0695 0.1000 0.1716
giou box_count_recall pixelwise lac multiplicative 14.4056 0.0262 24.2717 0.1000 35.0695 0.1000 0.1830
giou box_count_recall pixelwise aps additive 14.4056 0.0262 25.1563 0.0952 34.8062 0.0953 0.1602
giou box_count_recall pixelwise aps multiplicative 14.4056 0.0262 23.9725 0.0984 35.1776 0.0968 0.1805
giou box_count_recall boxwise lac additive 14.4056 0.0262 28.6734 0.0952 35.4345 0.0992 0.1716
giou box_count_recall boxwise lac multiplicative 14.4056 0.0262 25.0144 0.1000 35.4345 0.0992 0.1830
giou box_count_recall boxwise aps additive 14.4056 0.0262 28.6734 0.0952 34.8062 0.0953 0.1602
giou box_count_recall boxwise aps multiplicative 14.4056 0.0262 24.8396 0.0984 35.1776 0.0968 0.1805
hausdorff box_count_threshold thresholded lac additive 23.9744 0.0272 1.4441 0.0960 24.3973 0.1000 0.1728
hausdorff box_count_threshold thresholded lac multiplicative 23.9744 0.0272 1.0064 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold thresholded aps additive 23.9744 0.0272 1.4441 0.0924 23.4660 0.0984 0.1602
hausdorff box_count_threshold thresholded aps multiplicative 23.9744 0.0272 1.0064 0.0984 23.4660 0.0984 0.1797
hausdorff box_count_threshold pixelwise lac additive 23.9744 0.0272 1.0000 0.0984 24.3973 0.1000 0.1736
hausdorff box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold pixelwise aps additive 23.9744 0.0272 1.0000 0.0984 23.4660 0.0984 0.1602
hausdorff box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_threshold boxwise lac additive 23.9744 0.0272 1.3218 0.1048 24.3973 0.1000 0.1936
hausdorff box_count_threshold boxwise lac multiplicative 23.9744 0.0272 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_threshold boxwise aps additive 23.9744 0.0272 1.3218 0.1048 23.4660 0.0984 0.1797
hausdorff box_count_threshold boxwise aps multiplicative 23.9744 0.0272 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_recall thresholded lac additive 14.4056 0.0262 1.4441 0.0960 24.3973 0.1000 0.1728
hausdorff box_count_recall thresholded lac multiplicative 14.4056 0.0262 1.0064 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall thresholded aps additive 14.4056 0.0262 1.4441 0.0924 23.4660 0.0984 0.1602
hausdorff box_count_recall thresholded aps multiplicative 14.4056 0.0262 1.0064 0.0984 23.4660 0.0984 0.1797
hausdorff box_count_recall pixelwise lac additive 14.4056 0.0262 1.0000 0.0984 24.3973 0.1000 0.1736
hausdorff box_count_recall pixelwise lac multiplicative 14.4056 0.0262 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall pixelwise aps additive 14.4056 0.0262 1.0000 0.0984 23.4660 0.0984 0.1602
hausdorff box_count_recall pixelwise aps multiplicative 14.4056 0.0262 1.0000 0.1016 23.4660 0.0984 0.1797
hausdorff box_count_recall boxwise lac additive 14.4056 0.0262 1.3218 0.1048 24.3973 0.1000 0.1936
hausdorff box_count_recall boxwise lac multiplicative 14.4056 0.0262 1.0000 0.1016 24.3973 0.1000 0.1830
hausdorff box_count_recall boxwise aps additive 14.4056 0.0262 1.3218 0.1048 23.4660 0.0984 0.1797
hausdorff box_count_recall boxwise aps multiplicative 14.4056 0.0262 1.0000 0.1016 23.4660 0.0984 0.1797
lac box_count_threshold thresholded lac additive 23.9744 0.0272 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_threshold thresholded lac multiplicative 23.9744 0.0272 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_threshold thresholded aps additive 23.9744 0.0272 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_threshold thresholded aps multiplicative 23.9744 0.0272 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_threshold pixelwise lac additive 23.9744 0.0272 10.9243 0.1000 0.6689 0.1008 0.1830
lac box_count_threshold pixelwise lac multiplicative 23.9744 0.0272 6.6930 0.1016 0.6689 0.1008 0.1830
lac box_count_threshold pixelwise aps additive 23.9744 0.0272 10.9243 0.1000 0.8726 0.0780 0.1582
lac box_count_threshold pixelwise aps multiplicative 23.9744 0.0272 6.6930 0.1016 0.8726 0.0780 0.1582
lac box_count_threshold boxwise lac additive 23.9744 0.0272 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_threshold boxwise lac multiplicative 23.9744 0.0272 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_threshold boxwise aps additive 23.9744 0.0272 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_threshold boxwise aps multiplicative 23.9744 0.0272 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_recall thresholded lac additive 14.4056 0.0262 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_recall thresholded lac multiplicative 14.4056 0.0262 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_recall thresholded aps additive 14.4056 0.0262 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_recall thresholded aps multiplicative 14.4056 0.0262 8.3136 0.1000 0.8726 0.0780 0.1582
lac box_count_recall pixelwise lac additive 14.4056 0.0262 10.9243 0.1000 0.6689 0.1008 0.1830
lac box_count_recall pixelwise lac multiplicative 14.4056 0.0262 6.6930 0.1016 0.6689 0.1008 0.1830
lac box_count_recall pixelwise aps additive 14.4056 0.0262 10.9243 0.1000 0.8726 0.0780 0.1582
lac box_count_recall pixelwise aps multiplicative 14.4056 0.0262 6.6930 0.1016 0.8726 0.0780 0.1582
lac box_count_recall boxwise lac additive 14.4056 0.0262 12.2759 0.0952 0.6689 0.1008 0.1676
lac box_count_recall boxwise lac multiplicative 14.4056 0.0262 8.3136 0.1000 0.6689 0.1008 0.1797
lac box_count_recall boxwise aps additive 14.4056 0.0262 12.2759 0.0952 0.8726 0.0780 0.1516
lac box_count_recall boxwise aps multiplicative 14.4056 0.0262 8.3136 0.1000 0.8726 0.0780 0.1582
TABLE VII: Experimental Results using the YOLOv8x model for $\alpha^{\mathrm{tot}}=0.1$
Matching Function Conf. Loss Loc. Loss Cls. Pred. Set Loc. Pred. Set. Confidence Localization Classification Global
Size Risk Size Risk Size Risk Risk
mix box_count_threshold thresholded lac additive 17.7521 0.0179 10.9197 0.0484 12.5311 0.0476 0.0846
mix box_count_threshold thresholded lac multiplicative 17.7521 0.0179 6.4985 0.0512 12.5311 0.0476 0.0915
mix box_count_threshold thresholded aps additive 17.7521 0.0179 10.9197 0.0484 15.0931 0.0460 0.0857
mix box_count_threshold thresholded aps multiplicative 17.7521 0.0179 6.0228 0.0512 12.0499 0.0484 0.0923
mix box_count_threshold pixelwise lac additive 17.7521 0.0179 4.6034 0.0499 1.0064 0.0511 0.0871
mix box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 2.8759 0.0512 1.0064 0.0511 0.0915
mix box_count_threshold pixelwise aps additive 17.7521 0.0179 4.6034 0.0499 1.0000 0.0491 0.0868
mix box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 2.8759 0.0512 1.0000 0.0491 0.0915
mix box_count_threshold boxwise lac additive 17.7521 0.0179 5.1730 0.0499 1.0064 0.0511 0.0871
mix box_count_threshold boxwise lac multiplicative 17.7521 0.0179 3.6850 0.0512 1.0064 0.0511 0.0915
mix box_count_threshold boxwise aps additive 17.7521 0.0179 5.1730 0.0499 1.0000 0.0491 0.0868
mix box_count_threshold boxwise aps multiplicative 17.7521 0.0179 3.6850 0.0512 1.0000 0.0491 0.0915
mix box_count_recall thresholded lac additive 10.6552 0.0186 10.9197 0.0484 12.5311 0.0476 0.0846
mix box_count_recall thresholded lac multiplicative 10.6552 0.0186 6.4985 0.0512 12.5311 0.0476 0.0915
mix box_count_recall thresholded aps additive 10.6552 0.0186 10.9197 0.0484 15.0931 0.0460 0.0857
mix box_count_recall thresholded aps multiplicative 10.6552 0.0186 6.0228 0.0512 12.0499 0.0484 0.0923
mix box_count_recall pixelwise lac additive 10.6552 0.0186 4.6034 0.0499 1.0064 0.0511 0.0871
mix box_count_recall pixelwise lac multiplicative 10.6552 0.0186 2.8759 0.0512 1.0064 0.0511 0.0915
mix box_count_recall pixelwise aps additive 10.6552 0.0186 4.6034 0.0499 1.0000 0.0491 0.0868
mix box_count_recall pixelwise aps multiplicative 10.6552 0.0186 2.8759 0.0512 1.0000 0.0491 0.0915
mix box_count_recall boxwise lac additive 10.6552 0.0186 5.1730 0.0499 1.0064 0.0511 0.0871
mix box_count_recall boxwise lac multiplicative 10.6552 0.0186 3.6850 0.0512 1.0064 0.0511 0.0915
mix box_count_recall boxwise aps additive 10.6552 0.0186 5.1730 0.0499 1.0000 0.0491 0.0868
mix box_count_recall boxwise aps multiplicative 10.6552 0.0186 3.6850 0.0512 1.0000 0.0491 0.0915
giou box_count_threshold thresholded lac additive 17.7521 0.0179 26.2773 0.0476 35.0525 0.0484 0.0860
giou box_count_threshold thresholded lac multiplicative 17.7521 0.0179 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_threshold thresholded aps additive 17.7521 0.0179 26.2773 0.0476 37.3819 0.0460 0.0853
giou box_count_threshold thresholded aps multiplicative 17.7521 0.0179 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_threshold pixelwise lac additive 17.7521 0.0179 22.4286 0.0491 35.0525 0.0484 0.0868
giou box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 19.1563 0.0509 35.0525 0.0484 0.0909
giou box_count_threshold pixelwise aps additive 17.7521 0.0179 22.4286 0.0491 37.3819 0.0460 0.0853
giou box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 18.9911 0.0509 31.2916 0.0484 0.0909
giou box_count_threshold boxwise lac additive 17.7521 0.0179 23.3271 0.0491 35.0525 0.0484 0.0868
giou box_count_threshold boxwise lac multiplicative 17.7521 0.0179 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_threshold boxwise aps additive 17.7521 0.0179 23.3271 0.0491 37.3819 0.0460 0.0853
giou box_count_threshold boxwise aps multiplicative 17.7521 0.0179 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_recall thresholded lac additive 10.6552 0.0186 26.2773 0.0476 35.0525 0.0484 0.0860
giou box_count_recall thresholded lac multiplicative 10.6552 0.0186 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_recall thresholded aps additive 10.6552 0.0186 26.2773 0.0476 37.3819 0.0460 0.0853
giou box_count_recall thresholded aps multiplicative 10.6552 0.0186 19.6941 0.0499 31.2916 0.0484 0.0909
giou box_count_recall pixelwise lac additive 10.6552 0.0186 22.4286 0.0491 35.0525 0.0484 0.0868
giou box_count_recall pixelwise lac multiplicative 10.6552 0.0186 19.1563 0.0509 35.0525 0.0484 0.0909
giou box_count_recall pixelwise aps additive 10.6552 0.0186 22.4286 0.0491 37.3819 0.0460 0.0853
giou box_count_recall pixelwise aps multiplicative 10.6552 0.0186 18.9911 0.0509 31.2916 0.0484 0.0909
giou box_count_recall boxwise lac additive 10.6552 0.0186 23.3271 0.0491 35.0525 0.0484 0.0868
giou box_count_recall boxwise lac multiplicative 10.6552 0.0186 20.1749 0.0499 35.0525 0.0484 0.0909
giou box_count_recall boxwise aps additive 10.6552 0.0186 23.3271 0.0491 37.3819 0.0460 0.0853
giou box_count_recall boxwise aps multiplicative 10.6552 0.0186 19.6941 0.0499 31.2916 0.0484 0.0909
hausdorff box_count_threshold thresholded lac additive 17.7521 0.0179 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_threshold thresholded lac multiplicative 17.7521 0.0179 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold thresholded aps additive 17.7521 0.0179 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_threshold thresholded aps multiplicative 17.7521 0.0179 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_threshold pixelwise lac additive 17.7521 0.0179 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 2.5616 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold pixelwise aps additive 17.7521 0.0179 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 2.5616 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_threshold boxwise lac additive 17.7521 0.0179 3.9352 0.0510 34.3152 0.0491 0.0887
hausdorff box_count_threshold boxwise lac multiplicative 17.7521 0.0179 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_threshold boxwise aps additive 17.7521 0.0179 3.9352 0.0510 32.8959 0.0476 0.0879
hausdorff box_count_threshold boxwise aps multiplicative 17.7521 0.0179 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall thresholded lac additive 10.6552 0.0186 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_recall thresholded lac multiplicative 10.6552 0.0186 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall thresholded aps additive 10.6552 0.0186 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_recall thresholded aps multiplicative 10.6552 0.0186 2.9569 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall pixelwise lac additive 10.6552 0.0186 4.6034 0.0499 34.3152 0.0491 0.0883
hausdorff box_count_recall pixelwise lac multiplicative 10.6552 0.0186 2.5616 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall pixelwise aps additive 10.6552 0.0186 4.6034 0.0499 32.8959 0.0476 0.0879
hausdorff box_count_recall pixelwise aps multiplicative 10.6552 0.0186 2.5616 0.0512 32.8959 0.0476 0.0915
hausdorff box_count_recall boxwise lac additive 10.6552 0.0186 3.9352 0.0510 34.3152 0.0491 0.0887
hausdorff box_count_recall boxwise lac multiplicative 10.6552 0.0186 2.9569 0.0512 34.3152 0.0491 0.0915
hausdorff box_count_recall boxwise aps additive 10.6552 0.0186 3.9352 0.0510 32.8959 0.0476 0.0879
hausdorff box_count_recall boxwise aps multiplicative 10.6552 0.0186 2.9569 0.0512 32.8959 0.0476 0.0915
lac box_count_threshold thresholded lac additive 17.7521 0.0179 18.2558 0.0491 1.0064 0.0511 0.0868
lac box_count_threshold thresholded lac multiplicative 17.7521 0.0179 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_threshold thresholded aps additive 17.7521 0.0179 18.2558 0.0491 1.0000 0.0491 0.0864
lac box_count_threshold thresholded aps multiplicative 17.7521 0.0179 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_threshold pixelwise lac additive 17.7521 0.0179 14.4720 0.0499 1.0064 0.0511 0.0871
lac box_count_threshold pixelwise lac multiplicative 17.7521 0.0179 10.8107 0.0512 1.0064 0.0511 0.0915
lac box_count_threshold pixelwise aps additive 17.7521 0.0179 14.4720 0.0499 1.0000 0.0491 0.0868
lac box_count_threshold pixelwise aps multiplicative 17.7521 0.0179 10.8107 0.0512 1.0000 0.0491 0.0915
lac box_count_threshold boxwise lac additive 17.7521 0.0179 15.8677 0.0499 1.0064 0.0511 0.0871
lac box_count_threshold boxwise lac multiplicative 17.7521 0.0179 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_threshold boxwise aps additive 17.7521 0.0179 15.8677 0.0499 1.0000 0.0491 0.0868
lac box_count_threshold boxwise aps multiplicative 17.7521 0.0179 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_recall thresholded lac additive 10.6552 0.0186 18.2558 0.0491 1.0064 0.0511 0.0868
lac box_count_recall thresholded lac multiplicative 10.6552 0.0186 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_recall thresholded aps additive 10.6552 0.0186 18.2558 0.0491 1.0000 0.0491 0.0864
lac box_count_recall thresholded aps multiplicative 10.6552 0.0186 12.1580 0.0509 1.0000 0.0491 0.0909
lac box_count_recall pixelwise lac additive 10.6552 0.0186 14.4720 0.0499 1.0064 0.0511 0.0871
lac box_count_recall pixelwise lac multiplicative 10.6552 0.0186 10.8107 0.0512 1.0064 0.0511 0.0915
lac box_count_recall pixelwise aps additive 10.6552 0.0186 14.4720 0.0499 1.0000 0.0491 0.0868
lac box_count_recall pixelwise aps multiplicative 10.6552 0.0186 10.8107 0.0512 1.0000 0.0491 0.0915
lac box_count_recall boxwise lac additive 10.6552 0.0186 15.8677 0.0499 1.0064 0.0511 0.0871
lac box_count_recall boxwise lac multiplicative 10.6552 0.0186 12.1580 0.0509 1.0064 0.0511 0.0909
lac box_count_recall boxwise aps additive 10.6552 0.0186 15.8677 0.0499 1.0000 0.0491 0.0868
lac box_count_recall boxwise aps multiplicative 10.6552 0.0186 12.1580 0.0509 1.0000 0.0491 0.0909
TABLE VIII: Experimental Results using the YOLOv8x model for $\alpha^{\mathrm{tot}}=0.2$
Matching Function Conf. Loss Loc. Loss Cls. Pred. Set Loc. Pred. Set. Confidence Localization Classification Global
Size Risk Size Risk Size Risk Risk
mix box_count_threshold thresholded lac additive 15.4120 0.0219 7.1302 0.0932 10.0158 0.0960 0.1645
mix box_count_threshold thresholded lac multiplicative 15.4120 0.0219 4.1099 0.1008 10.0158 0.0960 0.1797
mix box_count_threshold thresholded aps additive 15.4120 0.0219 7.1302 0.0932 10.2455 0.0945 0.1645
mix box_count_threshold thresholded aps multiplicative 15.4120 0.0219 3.8031 0.1008 8.4829 0.0960 0.1797
mix box_count_threshold pixelwise lac additive 15.4120 0.0219 2.9879 0.0992 0.5591 0.1045 0.1797
mix box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 1.7427 0.1024 0.5591 0.1045 0.1830
mix box_count_threshold pixelwise aps additive 15.4120 0.0219 2.9879 0.0992 0.9931 0.0611 0.1449
mix box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 1.7427 0.1024 0.9817 0.0611 0.1473
mix box_count_threshold boxwise lac additive 15.4120 0.0219 3.4491 0.0992 0.5591 0.1045 0.1797
mix box_count_threshold boxwise lac multiplicative 15.4120 0.0219 2.2983 0.1024 0.5591 0.1045 0.1830
mix box_count_threshold boxwise aps additive 15.4120 0.0219 3.4491 0.0992 0.9931 0.0611 0.1449
mix box_count_threshold boxwise aps multiplicative 15.4120 0.0219 2.2983 0.1024 0.9817 0.0611 0.1473
mix box_count_recall thresholded lac additive 10.1384 0.0310 7.1302 0.0932 10.0158 0.0960 0.1645
mix box_count_recall thresholded lac multiplicative 10.1384 0.0310 4.1099 0.1008 10.0158 0.0960 0.1797
mix box_count_recall thresholded aps additive 10.1384 0.0310 7.1302 0.0932 10.2455 0.0945 0.1645
mix box_count_recall thresholded aps multiplicative 10.1384 0.0310 3.8031 0.1008 8.4829 0.0960 0.1797
mix box_count_recall pixelwise lac additive 10.1384 0.0310 2.9879 0.0992 0.5591 0.1045 0.1797
mix box_count_recall pixelwise lac multiplicative 10.1384 0.0310 1.7427 0.1024 0.5591 0.1045 0.1830
mix box_count_recall pixelwise aps additive 10.1384 0.0310 2.9879 0.0992 0.9931 0.0611 0.1449
mix box_count_recall pixelwise aps multiplicative 10.1384 0.0310 1.7427 0.1024 0.9817 0.0611 0.1473
mix box_count_recall boxwise lac additive 10.1384 0.0310 3.4491 0.0992 0.5591 0.1045 0.1797
mix box_count_recall boxwise lac multiplicative 10.1384 0.0310 2.2983 0.1024 0.5591 0.1045 0.1830
mix box_count_recall boxwise aps additive 10.1384 0.0310 3.4491 0.0992 0.9931 0.0611 0.1449
mix box_count_recall boxwise aps multiplicative 10.1384 0.0310 2.2983 0.1024 0.9817 0.0611 0.1473
giou box_count_threshold thresholded lac additive 15.4120 0.0219 19.5331 0.0960 26.8823 0.0968 0.1684
giou box_count_threshold thresholded lac multiplicative 15.4120 0.0219 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_threshold thresholded aps additive 15.4120 0.0219 19.5331 0.0960 26.6024 0.0953 0.1676
giou box_count_threshold thresholded aps multiplicative 15.4120 0.0219 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_threshold pixelwise lac additive 15.4120 0.0219 16.3220 0.0984 26.8823 0.0968 0.1712
giou box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 13.3474 0.1008 26.8823 0.0968 0.1797
giou box_count_threshold pixelwise aps additive 15.4120 0.0219 16.3220 0.0984 26.6024 0.0953 0.1676
giou box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 13.1761 0.1008 22.8826 0.0968 0.1797
giou box_count_threshold boxwise lac additive 15.4120 0.0219 17.0479 0.0984 26.8823 0.0968 0.1712
giou box_count_threshold boxwise lac multiplicative 15.4120 0.0219 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_threshold boxwise aps additive 15.4120 0.0219 17.0479 0.0984 26.6024 0.0953 0.1676
giou box_count_threshold boxwise aps multiplicative 15.4120 0.0219 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_recall thresholded lac additive 10.1384 0.0310 19.5331 0.0960 26.8823 0.0968 0.1684
giou box_count_recall thresholded lac multiplicative 10.1384 0.0310 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_recall thresholded aps additive 10.1384 0.0310 19.5331 0.0960 26.6024 0.0953 0.1676
giou box_count_recall thresholded aps multiplicative 10.1384 0.0310 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_recall pixelwise lac additive 10.1384 0.0310 16.3220 0.0984 26.8823 0.0968 0.1712
giou box_count_recall pixelwise lac multiplicative 10.1384 0.0310 13.3474 0.1008 26.8823 0.0968 0.1797
giou box_count_recall pixelwise aps additive 10.1384 0.0310 16.3220 0.0984 26.6024 0.0953 0.1676
giou box_count_recall pixelwise aps multiplicative 10.1384 0.0310 13.1761 0.1008 22.8826 0.0968 0.1797
giou box_count_recall boxwise lac additive 10.1384 0.0310 17.0479 0.0984 26.8823 0.0968 0.1712
giou box_count_recall boxwise lac multiplicative 10.1384 0.0310 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_recall boxwise aps additive 10.1384 0.0310 17.0479 0.0984 26.6024 0.0953 0.1676
giou box_count_recall boxwise aps multiplicative 10.1384 0.0310 13.7640 0.1000 22.8826 0.0968 0.1797
hausdorff box_count_threshold thresholded lac additive 15.4120 0.0219 4.7787 0.0901 25.4892 0.0984 0.1657
hausdorff box_count_threshold thresholded lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold thresholded aps additive 15.4120 0.0219 4.7787 0.0901 24.3324 0.0960 0.1657
hausdorff box_count_threshold thresholded aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_threshold pixelwise lac additive 15.4120 0.0219 2.9569 0.1008 25.4892 0.0984 0.1797
hausdorff box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold pixelwise aps additive 15.4120 0.0219 2.9569 0.1008 24.3324 0.0960 0.1765
hausdorff box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_threshold boxwise lac additive 15.4120 0.0219 3.9352 0.0960 25.4892 0.0984 0.1692
hausdorff box_count_threshold boxwise lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold boxwise aps additive 15.4120 0.0219 3.9352 0.0960 24.3324 0.0960 0.1692
hausdorff box_count_threshold boxwise aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall thresholded lac additive 10.1384 0.0310 4.7787 0.0901 25.4892 0.0984 0.1657
hausdorff box_count_recall thresholded lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall thresholded aps additive 10.1384 0.0310 4.7787 0.0901 24.3324 0.0960 0.1657
hausdorff box_count_recall thresholded aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall pixelwise lac additive 10.1384 0.0310 2.9569 0.1008 25.4892 0.0984 0.1797
hausdorff box_count_recall pixelwise lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall pixelwise aps additive 10.1384 0.0310 2.9569 0.1008 24.3324 0.0960 0.1765
hausdorff box_count_recall pixelwise aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall boxwise lac additive 10.1384 0.0310 3.9352 0.0960 25.4892 0.0984 0.1692
hausdorff box_count_recall boxwise lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall boxwise aps additive 10.1384 0.0310 3.9352 0.0960 24.3324 0.0960 0.1692
hausdorff box_count_recall boxwise aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
lac box_count_threshold thresholded lac additive 15.4120 0.0219 12.7995 0.0984 0.5591 0.1045 0.1757
lac box_count_threshold thresholded lac multiplicative 15.4120 0.0219 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_threshold thresholded aps additive 15.4120 0.0219 12.7995 0.0984 0.9931 0.0611 0.1449
lac box_count_threshold thresholded aps multiplicative 15.4120 0.0219 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_threshold pixelwise lac additive 15.4120 0.0219 9.6219 0.1000 0.5591 0.1045 0.1797
lac box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 6.8758 0.1016 0.5591 0.1045 0.1830
lac box_count_threshold pixelwise aps additive 15.4120 0.0219 9.6219 0.1000 0.9931 0.0611 0.1473
lac box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 6.8758 0.1016 0.9817 0.0611 0.1473
lac box_count_threshold boxwise lac additive 15.4120 0.0219 10.6310 0.1000 0.5591 0.1045 0.1797
lac box_count_threshold boxwise lac multiplicative 15.4120 0.0219 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_threshold boxwise aps additive 15.4120 0.0219 10.6310 0.1000 0.9931 0.0611 0.1473
lac box_count_threshold boxwise aps multiplicative 15.4120 0.0219 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_recall thresholded lac additive 10.1384 0.0310 12.7995 0.0984 0.5591 0.1045 0.1757
lac box_count_recall thresholded lac multiplicative 10.1384 0.0310 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_recall thresholded aps additive 10.1384 0.0310 12.7995 0.0984 0.9931 0.0611 0.1449
lac box_count_recall thresholded aps multiplicative 10.1384 0.0310 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_recall pixelwise lac additive 10.1384 0.0310 9.6219 0.1000 0.5591 0.1045 0.1797
lac box_count_recall pixelwise lac multiplicative 10.1384 0.0310 6.8758 0.1016 0.5591 0.1045 0.1830
lac box_count_recall pixelwise aps additive 10.1384 0.0310 9.6219 0.1000 0.9931 0.0611 0.1473
lac box_count_recall pixelwise aps multiplicative 10.1384 0.0310 6.8758 0.1016 0.9817 0.0611 0.1473
lac box_count_recall boxwise lac additive 10.1384 0.0310 10.6310 0.1000 0.5591 0.1045 0.1797
lac box_count_recall boxwise lac multiplicative 10.1384 0.0310 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_recall boxwise aps additive 10.1384 0.0310 10.6310 0.1000 0.9931 0.0611 0.1473
lac box_count_recall boxwise aps multiplicative 10.1384 0.0310 8.1480 0.1008 0.9817 0.0611 0.1473